### v1

# Schatten properties of commutators of fractional integrals on spaces of homogeneous type

Abstract.
Extending classical results of Janson and Peetre (1988) on the Schatten class $S^{p}$ membership of commutators of Riesz potentials on the Euclidean space, we obtain analogous results for commutators $[b,T]$, where $T\in\{T_{\varepsilon},\widetilde{T}_{\alpha}\}$ belongs to either one of two natural classes of fractional integral operators on a space of homogeneous type. Our approach is based on recent related work of Hytönen and Korte on singular (instead of fractional) integrals; working directly with the kernels, it differs from the Fourier analytic considerations of Janson and Peetre, covering new operators even when specialised to $\mathbb{R}^{d}$.
The cleanest case of our characterization in spaces of lower dimension $d>2$ and satisfying a $(1,2)$-Poincaré inequality is as follows. For a parameter $\varepsilon\in(0,\frac{1}{2}-\frac{1}{d})$ describing the order of the fractional integral $T_{\varepsilon}$, we have a dichotomy: If $\frac{d}{1+d\varepsilon}<p<\frac{1}{\varepsilon}$, then $b,T_{\varepsilon}]\in S^{p}$ if and only if $b$ belongs to a suitable Besov (or fractional Sobolev) space. If $0<p\leq\frac{d}{1+d\varepsilon}$, then $b,T_{\varepsilon}]\in S^{p}$ if and only if $b$ is constant. This is analogous to the result for singular integrals, where a similar cut-off happens at $p=d$, formally corresponding to fractional order $\varepsilon=0$. We also obtain results for other parameter values, including dimensions $0<d\leq 2$.
As an application, these results are used to show Schatten properties of commutators of fractional Bessel operators, complementing recent related results of Fan, Lacey, Li, and Xiong (2025) on commutators of singular integrals in the Bessel setting.
Key words and phrases:
Commutator, fractional integral, Schatten class, space of homogenous type, Besov space, fractional Sobolev space, Bessel operator, heat kernel
1991 Mathematics Subject Classification:
42B20, 42B35
T.H. was supported by the Research Council of Finland (project no. 364208). L.W. is supported by China Scholarship Council (Grant No. 202406310137)
Corresponding author.
1. Introduction
Fractional integral operators play a fundamental role in analysis due to their widespread applications in potential analysis, harmonic analysis, PDE and Sobolev embeddings.
The main classical example of fractional operators on the Euclidean space $\mathbb{R}^{d}$ with $d\geq 1$ is the Riesz potential $(-\Delta)^{-\frac{\alpha}{2}}$, with $\alpha=d\varepsilon>0$, given by
(1.1) $(-\Delta)^{-\frac{\alpha}{2}}f(x)=c_{d,\alpha}\int_{\mathbb{R}^{d}}\frac{f(y)}{|x-y|^{d-\alpha}}dy=c_{d,d\varepsilon}\int_{\mathbb{R}^{d}}\frac{f(y)}{|x-y|^{d(1-\varepsilon)}}dy,$
where
$c_{d,\alpha}=\frac{\Gamma(\frac{d-\alpha}{2})}{2^{\alpha}\pi^{\frac{d}{2}}\Gamma(\frac{\alpha}{2})};$
see [16, Chapter 6]. We write the two equivalent formulas above, parametrised by $\alpha$ and $\varepsilon$, since they give rise to two different classed of generalisations, as we will see below.
The topic of this paper is commutators of fractional integral operators with pointwise multipliers, namely, operators of the type
[b,T]f=bTfT(bf),[b,T]f=bTf-T(bf),
where $T=(-\Delta)^{-\frac{\alpha}{2}}$ or one of its generalisations that we shortly describe.
Some classical results in this theme are as follows: In [6], Chanillo showed that for any $0<\alpha<d$ and $1<p<\frac{d}{\alpha}$ as well as $\frac{1}{q}=\frac{1}{p}-\frac{\alpha}{d}$, the commutator $b,(-\Delta)^{-\frac{\alpha}{2}}]$ is bounded from $L^{p}(\mathbb{R}^{d})$ to $L^{q}(\mathbb{R}^{d})$ if and only if $b\in BMO(\mathbb{R}^{d})$, the space of functions of bounded mean oscillation. Later, Wang [35] proved that $b,(-\Delta)^{-\frac{\alpha}{2}}]$ is compact from $L^{p}(\mathbb{R}^{d})$ to $L^{q}(\mathbb{R}^{d})$ if and only if $b\in VMO(\mathbb{R}^{d})$, the $BMO$-closure of $C_{c}^{\infty}(\mathbb{R}^{d})$.
In this paper, we are particularly interested in quantitative versions of compactness measured in terms of the Schatten $S^{p}$ norms
(1.2) $\left\|R\right\|_{S^{p}(L^{2}(\mu))}:=\Big(\sum_{n=0}^{\infty}a_{n}(R)^{p}\Big)^{\frac{1}{p}},\qquad a_{n}(R):=\inf\{\left\|R-F\right\|_{L^{2}(\mu)\to L^{2}(\mu)}:\operatorname{rank}F\leq n\},$
where $a_{n}(R)$ is the $n$th approximation number (or singular value) of $R:L^{2}(\mu)\to L^{2}(\mu)$. In this direction, Janson and Peetre [26, p. 484] obtained the following results in the Euclidean space $\mathbb{R}^{d}$ with $d\geq 2$ as a special case of their work on so-called “paracommutators”:
  1. (1)
    For $p\geq 1$ and $\big(\frac{d}{p}-1\big)_{+}<\alpha<\min\{\frac{d}{p},\frac{d}{2}\}$, the commutator $b,{(-\Delta)}^{-\frac{\alpha}{2}}]$ belongs to the Schatten class $S^{p}(L^{2}(\mathbb{R}^{d}))$ if and only if $b$ belongs to the classical fractional Sobolev space $\dot{B}_{p,p}^{\frac{d}{p}-\alpha}(\mathbb{R}^{d})$ (see the definition in (2.13)).
  2. (2)
    For $1\leq p<d$ and $0<\alpha\leq\frac{d}{p}-1$, the commutator $b,{(-\Delta)}^{-\frac{\alpha}{2}}]$ belongs to the Schatten class $S^{p}(L^{2}(\mathbb{R}^{d}))$ if and only if $b$ is constant.
For $d\geq 2$ and $\varepsilon\in(0,\frac{1}{2})$, these results can be restated as follows:
(1.3) $b,(-\Delta)^{-\frac{d\varepsilon}{2}}]\in S^{p}(L^{2}(\mathbb{R}^{d}))\iff\begin{cases}b\in\dot{B}_{p,p}^{d(\frac{1}{p}-\varepsilon)}(\mathbb{R}^{d}),&\quad\frac{d}{1+d\varepsilon}<p<\frac{1}{\varepsilon},\\ b=\mathrm{constant},&\quad 0<p\leq\frac{d}{1+d\varepsilon}.\end{cases}$
At the critical point $p=\frac{d}{1+d\varepsilon}$, Frank, Sukochev, and Zanin [12] showed that $b,(-\Delta)^{-\frac{d\varepsilon}{2}}]$ belongs to the weak Schatten class $S^{p,\infty}$ if and only if $b$ belongs to the homogeneous Sobolev space $\dot{W}^{1}_{p}(\mathbb{R}^{d})$. For the corresponding Schatten $S^{p}$ properties of the fractional Laplacian operator $(-\Delta)^{\frac{\alpha}{2}}$ with positive order $\alpha>0$, we refer the reader to [12],[26],[30]. These questions fall outside the scope of the present work.
Several authors have also considered related questions in the more general setting of a space of homogeneous type $(X,\rho,\mu)$, which is a set $X$ with a quasi-distance $\rho$ and a positive measure $\mu$ such that the balls defined by $B(x,r)=\{y\in X:\rho(x,y)<r\}$ satisfy a doubling condition; see Section 2 for a detailed definition. In this setting, two types of fractional integrals are defined by
(1.4) $I_{\varepsilon}f(x)=\int_{X}\frac{1}{V(x,y)^{1-\varepsilon}}f(y)d\mu(y),\qquad\tilde{I}_{\alpha}f(x)=\int_{X}\frac{\rho(x,y)^{\alpha}}{V(x,y)}f(y)d\mu(y),$
where $V(x,y):=\mu(B(x,\rho(x,y)))$; we will refer to them as “volumic” and “metric”, respectively. When $X=\mathbb{R}^{d}$ with $\rho(x,y)=|x-y|$ and $d\mu=dx$, both $I_{\varepsilon}$ and $\tilde{I}_{\alpha}=\tilde{I}_{d\varepsilon}$ reduce to the classical Riesz potential $(-\Delta)^{-\frac{d\varepsilon}{2}}$. Note that $I_{\varepsilon}$ arises by interpreting the whole $|x-y|^{d(1-\varepsilon)}$ in (1.1) as $V(x,y)^{1-\varepsilon}$, while $\tilde{I}_{\alpha}$ is based on applying a different interpretation $|x-y|^{d}\sim V(x,y)$ and $|x-y|^{\alpha}=\rho(x,y)^{\alpha}$ to the two factors of $|x-y|^{d-\alpha}$.
Many works on fractional integrals over spaces of homogeneous type, like [14], are formulated in so-called “normal” spaces with $V(x,y)\sim\rho(x,y)$, in which case the volumic and the metric versions coincide. Without assuming normality, volumic fractional integrals have been studied e.g. in [31], and [5] obtained the $(L^{p},L^{q})$ boundedness of their commutators $b,I_{\varepsilon}]$ for $b\in BMO(X)$. These volumic fractional integrals admit a relatively clean theory in its own right. Nevertheless, it seems that metric fractional integrals, especially with $\alpha=1$, are actually the ones that more frequently arise in applications; see e.g. [19, Theorem 3.22 and Section 9.1] and [32, Eq. (7)], where further references to such operators in different contexts are given. In particular, fractional powers $(-\Delta_{\lambda})^{-\frac{\alpha}{2}}$ of the Bessel Laplacian $\Delta_{\lambda}$ turn out to be of the metric form; see [3] and Section 12 below. Incorporating this prominent example into our theory was a major motivation for dealing with $\tilde{I}_{\alpha}$, and we will return to this example in more detail below. More generally, we show in Section 11 that a large class of fractional operators arising from heat kernels fall under the umbrella of metric fractional integrals.
The aim of this paper is to investigate the Schatten class $S^{p}$ properties of the commutators $[b,T]$, where $T\in\{T_{\varepsilon},\widetilde{T_{\alpha}}\}$ belongs to one of two classes of fractional integral operators modelled after $I_{\varepsilon}$ and $\tilde{I}_{\alpha}$, respectively. We will show that $b,T]\in S^{p}$ if and only if $b$ belongs to a suitable Besov space, with certain fractional oscillatory spaces as intermediate steps in proving this equivalence.
Our approach is based on the recent works of Hytönen and Korte [23, 25] where, building on the work of Janson–Peetre [26] and Rochberg–Semmes [33] in the Euclidean case, they established similar results on spaces of homogeneous type for singular instead of fractional integrals, corresponding formally to the case $\varepsilon,\alpha=0$. Even in the Euclidean setting $X=\mathbb{R}^{d}$, our framework complements the results of Janson and Peetre [26], whose operators are defined on the Fourier transform side, in contrast to the more direct spatial description in our theory. While the basic case of $b,(-\Delta)^{-\frac{\alpha}{2}}]$ is covered by both, the examples beyond that are not comparable.
The following corollary, with clean conclusions under somewhat stronger assumptions than our main Theorem 2.14 and Corollary 2.15, serves as an illustration of our results.
Corollary 1.1.
Let $(X,\rho,\mu)$ be a space of homogeneous type with lower dimension $d>2$ (Definition 2.1) and satisfying the $(1,2)$-Poincaré inequality (Definition 2.12). Suppose that $\varepsilon\in(0,\frac{1}{2}-\frac{1}{d})$ and $\alpha=d\varepsilon\in(0,\frac{d}{2}-1)$. Then the following hold for all $b\in L_{\mathrm{loc}}^{1}(X)$:
  1. (i)
    If $p\in(\frac{d}{1+d\varepsilon},\frac{1}{\varepsilon})=(\frac{d}{1+\alpha},\frac{d}{\alpha})$, then {[b,Iε]Sp(L2(μ))  b ∈˙Bp1p(μ),[b,~Iα]Sp(L2(μ))  b ∈~Bpα(μ), where the two Besov spaces on the right are defined in (2.17) and (2.16), respectively.
  2. (ii)
    If $p\in(0,\frac{d}{1+d\varepsilon}]=(0,\frac{d}{1+\alpha}]$, then $b,I_{\varepsilon}]\in S^{p}(L^{2}(\mu))$ or $b,\tilde{I}_{\alpha}]\in S^{p}(L^{2}(\mu))$ if and only if $b$ is constant.
More generally, the same conclusions hold for all strongly non-degenerate $\phi$-fractional integral operators (Definition 2.6) $T_{\varepsilon}$ in place of $I_{\varepsilon}$ and $\widetilde{T}_{\alpha}$ in place of $\tilde{I}_{\alpha}$, where $\phi(x,y)\in\{V(x,y)^{\varepsilon},\rho(x,y)^{\alpha}\}$, respectively.
We note that the $(1,2)$-Poincaré inequality is a natural assumption, in the sense that it is the version of the Poincaré inequality most frequently established in concrete situations; see e.g. Baudoin et al. [1].
Remark 1.2.
For $(X,\rho,\mu)=(\mathbb{R}^{d},|x-y|,dx)$, both $\dot{B}_{p}^{\frac{1}{p}-\varepsilon}(\mu)$ and $\widetilde{B}_{p}^{d\varepsilon}(\mu)$ coincide with the classical Besov space $\dot{B}_{p,p}^{d(\frac{1}{p}-\varepsilon)}(\mathbb{R}^{d})$, and Corollary 1.1 applies in particular to $I_{\varepsilon}=\tilde{I}_{d\varepsilon}=(-\Delta)^{-\frac{d\varepsilon}{2}}$. We see that the Corollary reproduces the classical Janson–Peetre result (1.3), except for the fact that our restrictions on the dimension (our $d>2$ vs. $d\geq 2$ in (1.3)) and the fractional parameter (our $\varepsilon\in(0,\frac{1}{2}-\frac{1}{d})$ vs. $\varepsilon\in(0,\frac{1}{2})$ in (1.3)) are somewhat stronger. These restrictions arise from a limitation of our method described in Remark 2.16. On the other hand, in this smaller range, Corollary 1.1 not only recovers (1.3) for $(-\Delta)^{-\frac{\alpha}{2}}$, but also covers a large range of other fractional operators as in Definition 2.6.
As a more serious application of our abstract results, we characterise the Schatten properties of commutators of fractional powers of the Bessel Laplacian
(1.5) $\Delta_{\lambda}^{(n+1)}:=\frac{\partial^{2}}{\partial x_{1}^{2}}+\dots+\frac{\partial^{2}}{\partial x_{n}^{2}}+\frac{\partial^{2}}{\partial x_{n+1}^{2}}+\frac{2\lambda}{x_{n+1}}\cdot\frac{\partial}{\partial x_{n+1}}$
on $\mathbb{R}_{+}^{n+1}=\mathbb{R}^{n}\times(0,\infty)$. We give here the following illustrative result, leaving a more general statement for Corollary 12.1:
Corollary 1.3.
Let $n\geq 2$, $\lambda>0$, and $0<\alpha<\frac{n-1}{2}$. Let $(-\Delta_{\lambda})^{-\alpha/2}$ be the fractional Bessel operator in $(\mathbb{R}^{n+1}_{+},|\cdot|,dm_{\lambda}^{(n+1)})$, where $dm_{\lambda}^{(n+1)}(x)=x_{n+1}^{2\lambda}dx$. Then the following conclusions hold for all $b\in L_{\mathrm{loc}}^{1}(\mathbb{R}^{n+1}_{+})$:
$b,(-\Delta_{\lambda})^{-\frac{\alpha}{2}}]\in S^{p}(L^{2}(dm_{\lambda}^{(n+1)}))\quad\iff\quad\begin{cases}b\in\widetilde{B}_{p}^{\alpha}(dm_{\lambda}^{(n+1)}),&\text{if}\quad p\in(\frac{n+1}{\alpha+1},\frac{n+1}{\alpha}),\\ b=\text{const},&\text{if}\quad p\in(0,\frac{n+1}{\alpha+1}],\end{cases}$
where $\widetilde{B}_{p}^{\alpha}(dm_{\lambda}^{(n+1)})$ is defined as in (2.16) with $(X,\rho,\mu)=(\mathbb{R}^{n+1}_{+},|\cdot|,dm_{\lambda}^{(n+1)})$.
Sketch of proof.
This is a direct application of Corollary 1.1, once we verify the following:
  1. (i)
    $(\mathbb{R}^{n+1}_{+},|\cdot|,dm_{\lambda}^{(n+1)})$ is a space of homogeneous type of lower dimension $d=n+1>2$ that satisfies the $(1,2)$ (in fact, even the stronger $(1,1)$) Poincaré inequality. This is [23, Proposition 4.2], restated as Proposition 12.3.
  2. (ii)
    $(-\Delta_{\lambda})^{-\frac{\alpha}{2}}$ is a strongly non-degenerate $\phi$-fractional integral operator on this space with $\phi(x,y)=\left|x-y\right|^{\alpha}$. This is Proposition 12.5.
Part (ii) of the proof above is due to [3] in the special case $n=0$ (i.e., on $\mathbb{R}_{+}=(0,\infty)$), but seems to be unavailable in the previous literature for $n\geq 1$. We establish these properties in Section 12. Besides Corollary 1.3, this may have independent interest in bringing the fractional Bessel operators $(-\Delta_{\lambda})^{-\frac{\alpha}{2}}$ under the general umbrella of fractional integrals on spaces of homogeneous type, for which other results can then be directly quoted from the literature.
There are previous results related to Corollary 1.3 by Fan, Lacey, Li, and Xiong [11], who deal with the Schatten properties of commutators associated with the Bessel Riesz transforms $R_{\lambda,j}=\partial_{j}(-\Delta_{\lambda})^{-\frac{1}{2}}$. By the kernel estimates of $R_{\lambda,j}$ obtained in [10], these are operators of singular integral type, formally corresponding to $\widetilde{T}_{\alpha}$ with $\alpha=0$, which are in the scope of the theory of Hytönen and Korte [23, 25].
The structure of the paper is as follows. In Section 2, we give the main definitions and state the general form of our main results. In Section 3, we prove some basic lemmas supporting the definition in Section 2, and we introduce the definition of dyadic cubes as preparatory tools. Section 4 is devoted to establishing upper Schatten bounds of the commutators $[b,T]$. In Section 5, we present the equivalent characterization of the fractional Sobolev norms via some fractional oscillatory norms. In Section 6, we simplify the recent complex median method of Wei and Zhang [36] and apply it to fractional integrals. Building on the results from Section 5 and 6, we derive the lower Schatten bounds of the commutators in Section 7. Adapting results from Hytönen and Korte [25], we show in Section 8 that certain Besov spaces only consist of constants. The proofs the main results, including Corollary 1.1, are then completed by synthesizing the preceding estimates in Section 9.
The final three sections provide examples of fractional integrals that fall under the scope of our theory. In Section 10, we study kernels with additional regularity, which is often available in applications. In Section 11, we show that negative fractional powers $\mathcal{L}^{-s}$ of generators of heat semigroups $e^{-t\mathcal{L}}$, under quite general assumptions on the heat kernel $p_{t}(x,y)$, are metric fractional integrals in the sense of our definition. In Section 12, we deal with the specific case of fractional Bessel operators, after verifying that they fit into our general framework.
Notation
We write $X\lesssim Y$ to mean $X\leq CY$ for some constant $C>0$ independent of key variables, and $X\sim Y$ when both $X\lesssim Y$ and $Y\lesssim X$ hold.
2. Definitions and main results
We now provide the full set of relevant definitions and then state the general form of our main results, a special case of which was formulated in Corollary 1.1 in the Introduction.
Firstly, we recall that $(X,\rho,\mu)$ is a space of homogeneous type if $\rho:X\times X\rightarrow[0,+\infty)$ is a quasi-metric on the set $X$ satisfying the following properties: $\mathrm{(i)}$ $\rho(x,y)=\rho(y,x)\geq 0$ for all $x,y\in X$, $\mathrm{(ii)}$ $\rho(x,y)=0$ if and only if $x=y$, $\mathrm{(iii)}$ there exists a constant $A_{0}\geq 1$ such that for all $x,y,z\in X$,
(2.1) $\rho(x,y)\leq A_{0}[\rho(x,z)+\rho(z,y)],$
and $\mu$ is a positive Borel measure on $X$, satisfying the doubling condition:
(2.2) $0<\mu(2B)\leq C\mu(B)<\infty,\quad\text{for all balls }B.$
We abbreviate
$V(x,r):=\mu(B(x,r)),\qquad V(x,y):=\begin{cases}V(x,\rho(x,y)),&x\neq y,\\ \mu(\{x\}),&x=y.\end{cases}$
Definition 2.1.
We say that $(X,\rho,\mu)$ has upper dimension $D>0$ if there exists a constant $C_{\mu}\geq 1$ such that for any $x\in X$ and $0<r\leq R<\infty$,
(2.3) $\frac{V(x,R)}{V(x,r)}\leq C_{\mu}(\frac{R}{r})^{D},$
and $(X,\rho,\mu)$ has lower dimension $d>0$ if there exists a constant $\widetilde{C}_{\mu}\geq 1$ such that for any $x\in X$ and $0<r\leq R<\infty$,
(2.4) $\frac{V(x,R)}{V(x,r)}\geq\widetilde{C}_{\mu}(\frac{R}{r})^{d}.$
Note that (2.3) is equivalent to the doubling condition (2.2).
A measure $\mu$ on $X$ is said to be Ahlfors $\gamma$-regular, if there is a constant $\beta\geq 1$ such that
${\beta}^{-1}r^{\gamma}\leq\mu(B(x,r))\leq\beta r^{\gamma},$
for any $0<r<\infty$ and any ball $B(x,r)$ in $X$. A metric space $X$ carrying an Ahlfors $\gamma$-regular measure is called an Ahlfors $\gamma$-regular space. It has both the upper dimension $D=\gamma$ and the lower dimension $d=\gamma$.
2.1. Fractional integrals and their commutators
To streamline the discussion of the two different versions of fractional integrals, for $\varepsilon\in(0,1)$ and $\alpha\in(0,\infty)$, we denote
(2.5) $\phi(x,r)=\phi(B(x,r))\in\{V(x,r)^{\varepsilon},r^{\alpha}\}.$
and, consistently with the two cases in (2.5),
(2.6) $\phi(x,y):=\phi(x,\rho(x,y))\in\{V(x,y)^{\varepsilon},\rho(x,y)^{\alpha}\}.$
Definition 2.2.
A function $K\in L_{\mathrm{loc}}^{1}(X\times X)$ is called a $\phi$-fractional integral kernel if there is a constant $C_{K}$ such that
(2.7) $|K(x,y)|\leq C_{K}\frac{\phi(x,y)}{V(x,y)}\qquad\Big(\text{with}\quad\frac{0}{0}:=0\Big)$
for all $x,y\in X$.
Definition 2.3.
A $\phi$-fractional integral kernel $K$ is said to be strongly non-degenerate if there are positive constants $A,c_{1}$ and $\overline{C}$, and $\eta\leq\frac{\pi}{9}$, such that for every $x_{0}\in X$ and $r>0$, there exists a point $y_{0}\in B(x_{0},\overline{C}Ar)\setminus B(x,Ar)$ such that for some $v\in\mathbb{C}$ with $\left|v\right|=1$, we have at least one of the following two options:
(2.8) $|K(x,y)|\geq c_{1}\cdot\frac{\phi(x_{0},r)}{V(x_{0},r)},\qquad|\arg(\bar{v}K(x,y))|\leq\eta$
for all $x\in B(x_{0},r)$ and $y\in B(y_{0},r)$, or
(2.9) $|K(y,x)|\geq c_{1}\cdot\frac{\phi(x_{0},r)}{V(x_{0},r)},\qquad|\arg(\bar{v}K(y,x))|\leq\eta$
for all $x\in B(x_{0},r)$ and $y\in B(y_{0},r)$.
Remark 2.4.
Although “strongly non-degenerate” is the main notion of non-degeneracy that we use in this paper, we reserve the simpler name “non-degenerate” for another variant (Definition 10.1), since this variant is closer to the notion of “non-degenerate” for singular integral kernels as defined in [23, Eq. (1.7)].
The existence of a (strongly) non-degenerate kernel requires in particular that $B(x,r)\neq X$ for all $r>0$, hence that $X$ is unbounded, and thus (by [5, Lemma 1.9]) that $\mu(X)=\infty$. Hence, this assumption is implicitly in force in all results dealing with (strongly) non-degenerate kernels. A modification can be made to accommodate spaces of finite diameter; see [23, Section 10] for related discussion in the case of singular (instead of fractional) kernels.
Moreover, the existence of a (strongly) non-degenerate kernel also implies the non-empty annulus property $B(x,\overline{C}r)\setminus B(x,r)\neq\varnothing$ for all $x\in X$ and $r>0$, which is equivalent to inequality (2.4) for some $d>0$ by [18, Remark 1.2]. In particular, this implies that $\mu(\{x\})=0$ for all $x\in X$, which is hence also implicitly assumed in all results dealing with (strongly) non-degenerate kernels.
In many cases, strong non-degeneracy can be deduced from a simple two-sided bound for a non-negative fractional kernel. A prominent example is the fractional Bessel kernel, see (12.7). We will give the proof of Lemma 2.5 in Section 3.1.
Lemma 2.5.
Let $(X,\rho,\mu)$ be a space of homogeneous type with the non-empty annulus property $B(x,\overline{C}r)\setminus B(x,r)\neq\varnothing$ for all $x\in X$ and $r>0$. Let $K$ be a $\phi$-fractional integral kernel that satisfies
(2.10) $K(x,y)\sim\frac{\phi(x,y)}{V(x,y)}$
for all $x,y\in X$. Then $\phi$ is strongly non-degenerate.
Definition 2.6.
Let $K$ be a $\phi$-fractional integral kernel satisfying (2.6) and (2.7). The associated $\phi$-fractional integral is defined by
(2.11) $Tf(x)=\int_{X}K(x,y)f(y)d\mu(y)$
for all $f\in L^{1}_{\operatorname{loc}}(X)$ and $x\in X$ for which the integral (2.11) is well defined. The operator $T$ is called strongly non-degenerate whenever its kernel $K$ has the corresponding property. We write $T\in\{T_{\varepsilon},\widetilde{T}_{\alpha}\}$ according to the two cases in (2.6).
Example 2.7.
The basic fractional integrals $I_{\varepsilon}$ and $\tilde{I}_{\alpha}$ from (1.4) are strongly non-degenerate $\phi$-fractional integrals with $\phi(x,y)=V(x,y)^{\varepsilon}$ and $\phi(x,y)=\rho(x,y)^{\alpha}$, respectively.
Proof..
That $I_{\varepsilon}$ and $\tilde{I}_{\alpha}$ from (1.4) are $\phi$-fractional integrals with the respective $\phi$ is clear; indeed, their kernels are equal to the upper bound defining a $\phi$-fractional integral kernel. The strong non-degeneracy follows from Lemma 2.5. ∎
The following basic lemma, whose proof we postpone to Section 3.1, guarantees that (2.11) is well defined for a rather rich class of functions:
Lemma 2.8.
Let $\phi$ and $K$ satisfy (2.6) and (2.7). Then
  1. (i)
    If $f\in L^{1}(X)$ is boundedly supported, then (2.11) is well defined for a.e. $x\in X$.
  2. (ii)
    If $f\in L^{p}(X)$ and $g\in L^{p^{\prime}}(X)$ are boundedly supported, where $p\in[1,\infty]$ and $\frac{1}{p}+\frac{1}{p^{\prime}}=1$, then $\langle Tf,g\rangle=\int_{X}Tf(x)g(x)d\mu(x)$ is well defined, and
    (2.12) $|\langle Tf,g\rangle|\leq\int_{X}\int_{X}|K(x,y)f(y)g(x)|d\mu(y)d\mu(x)\lesssim\phi(x_{0},r)\left\|f\right\|_{L^{p}(X)}\left\|g\right\|_{L^{p^{\prime}}(X)}$
    if the supports of $f$ and $g$ are contained in $B(x_{0},r)$.
Remark 2.9.
If the kernel $K$ satisfies (2.7) and (10.1) with $\varepsilon,\alpha=0$, then $K$ is called a standard kernel and the corresponding operator $T$ as in (2.11) is called a Calderón-Zygmund operator. Giving a meaning to the integral (2.11) is much trickier in this case, and one usually only requires that the representation (2.11) is valid for $x$ outside the support of $f$.
Our main object of study is commutators associated to the fractional integral $T$:
Definition 2.10.
Let $b\in L_{\mathrm{loc}}^{1}(X)$. Let $T$ be a $\phi$-fractional integral operator as in Definition 2.6, with $\phi$-fractional integral kernel. The commutator $[b,T]$ is defined by
$b,T]f(x):=\int_{X}(b(x)-b(y))K(x,y)f(y)d\mu(y)$
[…]
Lemma 2.11.
  1. (i)
    If both $Tf(x)$ and $T(bf)(x)$ are well defined, then so is [b,T]f(x)[b,T]f(x)‚ and we have
    [b,T]f(x)=b(x)Tf(x)T(bf)(x).[b,T]f(x)=b(x)Tf(x)-T(bf)(x).
  2. (ii)
    Case (i) holds in particular if $f\in L^{\infty}(X)$ is boundedly supported.
  3. (iii)
    If both $f,g\in L^{\infty}(X)$ are boundedly supported, then $\langle[b,T]f,g\rangle=\langle Tf,bg\rangle-\langle T(bf),g\rangle$ is well defined.
Proof..
(i) is immediate from the definitions.
(ii): Under this assumption both $f,bf\in L^{1}(X)$ are boundedly supported, and the claim follows from Lemma 2.8(i).
(iii): This follows from Lemma 2.8(ii) applied to both boundedly supported pairs of functions $(f,bg)\in L^{\infty}(X)\times L^{1}(X)$ and $(bf,g)\in L^{1}(X)\times L^{\infty}(X)$ in place of $(f,g)$. ∎
In particular, Lemma 2.11 shows that, under the very general assumption $b\in L^{1}_{\operatorname{loc}}(X)$, the commutator $[b,T]$ is well defined on a class of test functions that is dense in $L^{2}(X)$. Hence, the question of extending $[b,T]$ to a bounded operator on $L^{2}(X)$ is equivalent to estimates on this dense test class. We are mainly interested in the stronger property that $[b,T]$ is not only bounded on $L^{2}(X)$ but belongs to the Schatten class $S^{p}=S^{p}(L^{2}(X))$ of certain compact operators on $L^{2}(X)$. Our main results will provide sufficient and necessary conditions for this in the following sense:
  1. (i)
    If $b$ belongs to a suitable subclass of $L^{1}_{\operatorname{loc}}(X)$, then $[b,T]f(x)$ is well-defined for all $f\in L^{2}(X)$ and a.e. $x\in X$, and the operator $[b,T]$ thus defined belongs to $S^{p}$.
  2. (ii)
    If $b\in L^{1}_{\operatorname{loc}}(X)$ and the operator $[b,T]$, first defined on boundedly supported $f\in L^{\infty}(X)$ only, has an extension to a bounded linear operator on $L^{2}(X)$ of class $S^{p}$, then $b$ belongs to a suitable subclass of $L^{1}_{\operatorname{loc}}(X)$.
2.2. Fractional Sobolev norms and the Poincaré inequality
Suppose that $1<p<\infty$ and $0<s<1$. The classical fractional Sobolev space $\dot{B}_{p,p}^{s}(\mathbb{R}^{d})$ is defined as all locally integrable function $b$ on the Euclidean space $\mathbb{R}^{d}$ such that
(2.13) $\|b\|_{\dot{B}_{p,p}^{s}(\mathbb{R}^{d})}=\left(\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|b(x)-b(y)|^{p}}{|x-y|^{d+sp}}dx\,dy\right)^{\frac{1}{p}}<\infty.$
For many purposes (see e.g. [15]), its relevant extension to spaces of homogeneous type is defined by
(2.14) $\|b\|_{\dot{B}_{p,p}^{s}(\mu)}=\left(\int_{X}\int_{X}\left(\frac{|b(x)-b(y)|}{\rho(x,y)^{s}}\right)^{p}\frac{d\mu(x)d\mu(y)}{V(x,y)}\right)^{\frac{1}{p}}<\infty,$
where the integrand involves both $\rho$ and $V$, i.e., the factor $|x-y|^{d}$ in (2.13) is interpreted as a volume $V(x,y)$, but the factor $|x-y|^{sp}$ as a distance $\rho(x,y)^{sp}$.
However, for the study of $\phi$-fractional integrals $T$, the following variant seems more natural, and this will be confirmed by its appearance in the characterizing conditions of Theorem 2.14 below. For $p\in(1,\infty)$ and $\phi$ as in (2.6), let
(2.15) $\|b\|_{\mathbb{B}_{p}(\phi,\mu)}:=\left(\int_{X}\int_{X}\frac{|b(x)-b(y)|^{p}}{V(x,y)^{2}}\phi(x,y)^{p}d\mu(x)d\mu(y)\right)^{\frac{1}{p}}.$
When $\phi=\rho^{\alpha}$, we also denote
(2.16) $\|b\|_{\widetilde{B}_{p}^{\alpha}(\mu)}:=\|b\|_{\mathbb{B}_{p}(\rho^{\alpha},\mu)}:=\left(\int_{X}\int_{X}\frac{|b(x)-b(y)|^{p}}{V(x,y)^{2}}\cdot\rho(x,y)^{p\alpha}d\mu(x)d\mu(y)\right)^{\frac{1}{p}}.$
However, in the volumic case, we adopt a different normalisation
(2.17) $\|b\|_{\dot{B}_{p}^{\varepsilon}(\mu)}:=\|b\|_{\mathbb{B}_{p}(V^{\frac{1}{p}-\varepsilon},\mu)}:=\left(\int_{X}\int_{X}\frac{|b(x)-b(y)|^{p}}{V(x,y)^{1+p\varepsilon}}d\mu(x)d\mu(y)\right)^{\frac{1}{p}},$
noting that $V^{p(\frac{1}{p}-\varepsilon)-2}=V^{1-\varepsilon p-2}=V^{-1-\varepsilon p}$. The motivation of this normalisation is that the parameter $\varepsilon$ in $\dot{B}_{p}^{\varepsilon}(\mu)$ plays a similar role as the classical smoothness parameter $s$ in (2.14). Notably, if $(X,\rho,\mu)$ is Ahlfors $d$-regular, then
$\dot{B}^{s}_{p,p}(\mu)=\dot{B}^{\frac{s}{d}}_{p}(\mu)=\widetilde{B}^{\frac{d}{p}-s}_{p}(\mu)\quad\text{if}\quad V\sim\rho^{d}.$
With this normalisation, the volumic Besov space that appears in our results about $S^{p}$ properties of commutators will be
$\|b\|_{\dot{B}_{p}^{\frac{1}{p}-\varepsilon}(\mu)}:=\|b\|_{\mathbb{B}_{p}(V^{\varepsilon},\mu)}:=\left(\int_{X}\int_{X}\frac{|b(x)-b(y)|^{p}}{V(x,y)^{2}}V(x,y)^{\varepsilon p}d\mu(x)d\mu(y)\right)^{\frac{1}{p}};$
except for the dimensional factor $d$, this is similar to the form of the classical results as in (1.3).
A similar space also featured in analogous results for singular integrals in [23]. The space denoted by $\dot{B}_{p}(\mu)$ in [23] corresponds to $\dot{B}_{p}^{\frac{1}{p}}(\mu)=\tilde{B}^{0}_{p}(\mu)=\mathbb{B}_{p}(1,\mu)$ in the present notation, taking $\phi\equiv 1$ in (2.15).
We also recall the Poincaré inequality, which plays a significant role in several aspect of analysis on metric spaces (see [19]), and our main result below is no exception.
Definition 2.12.
Let $s>1$. A space $(X,\rho,\mu)$ is said to satisfy the $(1,s)$-Poincaré inequality if $\rho$ is a metric (i.e., $A_{0}=1$ in (2.1)), and there exists $\lambda\geq 1$ and $c_{P}$ such that for every Lipschitz function $f$ on $X$, every $x\in X$ and $r>0$,
(2.18) $\fint_{B(x,r)}|f-\langle f\rangle_{B(x,r)}|d\mu\leq c_{P}\cdot r\cdot\left(\fint_{B(x,\lambda r)}(\mathrm{lip}f)^{s}d\mu\right)^{1/s},$
where the pointwise Lipschitz constant $\mathrm{lip}f$ is defined as
(2.19) $\mathrm{lip}f(x):=\underset{r\rightarrow 0}{\lim\inf}\,\,\,\underset{\rho(x,y)\leq r}{\sup}\frac{|f(x)-f(y)|}{r}.$
Remark 2.13.
On metric spaces, the abundance of Lipschitz functions makes the Poincaré inequality a useful and non-trivial condition. To accommodate quasi-metric spaces, we incorporate the assumption that $\rho$ is a metric into the definition of the Poincaré inequality. This allows us to use the phrase “let $X$ satisfy the Poincaré inequality” as a shorthand for “let $X$ be a metric space that satisfies the Poincaré inequality”.
For any $s>1$, if $X$ satisfies the $(1,s)$-Poincaré inequality, then $X$ satisfies the $(1,t)$-Poincaré inequality for every $t\geq s$. This is immediate from Hölder’s inequality.
If $X$ is a complete doubling space, then it also satisfies the $(1,t)$-Poincaré inequality for some $t<s$. This is a deeper theorem from [28]; we will not need it in the present work.
2.3. The main results
We are now ready to state our main result. We will write simply $S^{p}:=S^{p}(L^{2}(\mu))$, where the space $L^{2}(\mu)$ is understood from the context.
Theorem 2.14.
Let $(X,\rho,\mu)$ be a space of homogeneous type with a lower dimension $d>0$. Let $\phi$ be as in (2.6) with parameter $\varepsilon\in(0,1)$ or $\alpha\in(0,\infty)$. Suppose that $T$ is a $\phi$-fractional integral. Then the following conclusions hold for all $b\in L_{\mathrm{loc}}^{1}(X)$:
  1. (1)
    If $p\in[2,\infty)$, $\varepsilon\in(0,1)$, and $\alpha\in(0,\infty)$, then [b,T_ε] _S^p b _˙B_p^1p-ε(μ)   and   [b,~T_α] _S^p b _~B_p^α(μ).
  2. (2)
    If $p\in(1,\infty)$, $\varepsilon\in(0,\frac{1}{p})$, $\alpha=d\varepsilon$, and $K$ is strongly non-degenerate (Definition 2.3), then b _˙B_p^1p-ε(μ) [b,T_ε] _S^p   and   b _~B_p^α(μ) [b,~T_α] _S^p .
  3. (3)
    If $d\in(1,\infty)$, $\varepsilon\in(0,1-\frac{1}{d})$, $\alpha=d\varepsilon$, and $X$ satisfies the $(1,\frac{d}{1+d\varepsilon})$-Poincaré inequality, then
    (2.20) $\dot{B}_{\frac{d}{1+d\varepsilon}}^{\frac{1}{d}}(\mu)=\widetilde{B}_{\frac{d}{1+\alpha}}^{\alpha}(\mu)\equiv\{\mathrm{constants}\}.$
The first result (1) does not require the existence of lower dimension $d$.
Combining the upper and lower bounds, we can further present the following corollary:
Corollary 2.15.
Let $(X,\rho,\mu)$ be a space of homogeneous type with a lower dimension $d>0$. Let $\varepsilon\in(0,1)$ and $\alpha=d\varepsilon$. Suppose that $T$ is a strongly non-degenerate fractional integral operator with kernel $K$ satisfying (2.7) through (10.2). Then the following conclusions hold for all $b\in L_{\mathrm{loc}}^{1}(X)$:
  1. (1)
    If $p\in[2,\infty)$ and $\varepsilon\in(0,\frac{1}{p})$, then {[b,Tε]Sp    b ∈˙Bp1p(μ),[b,~Tα]Sp    b ∈~Bpα(μ).
  2. (2)
    If $p\in(1,2)$ and $\varepsilon\in(\max\{0,\frac{1}{p}-\frac{1}{d}\},\frac{1}{p})$, then {[b,Tε]Sp    b ∈˙Bp1p(μ),[b,~Tα]Sp    b ∈~Bpα(μ).
  3. (3)
    If $p\in(0,d)$ and $\varepsilon\in(0,1-\frac{1}{d})\cap(0,\frac{1}{p}-\frac{1}{d}]$ and $X$ satisfies the $(1,\frac{d}{1+d\varepsilon})$-Poincaré inequality, then $b,T_{\varepsilon}]\in S^{p}$ or $b,\widetilde{T_{\alpha}}]\in S^{p}$ if and only if $b$ is constant.
$\tfrac{1}{p}$$\varepsilon$$\varepsilon=\tfrac{1}{p}$$\varepsilon=\tfrac{1}{p}-\tfrac{1}{d}$$\tfrac{1}{d}$$\tfrac{1}{2}$$\tfrac{1}{2}1$$\tfrac{1}{d}1$(1)(2)(3)
(a) Case: $0<d\leq 2$
[…]
(b) Case: $d>2$
Figure 1. The different parameter ranges for $(\varepsilon,p,d)$ in Corollary 2.15. In cases (1) and (3), we have a characterization, and in (2), a necessary condition for $b,T_{\varepsilon}]\in S^{p}$ and $b,\widetilde{T_{\alpha}}]\in S^{p}$ (with $\alpha=d\varepsilon$). The region below the dashed line $\varepsilon=\tfrac{1}{2}-\tfrac{1}{d}$ in case $d>2$ corresponds to the parameter ranges in Corollary 1.1, where only cases (1) and (3) appear.
Remark 2.16.
We point out that the method of estimating the upper Schatten bounds of the fractional commutator $[b,T]$ in Corollary 4.2 is limited to the parameter $p\in[2,\infty)$. The reverse of Corollary 2.15 (2) for $p\in(1,2)$ will be addressed in a forthcoming work of the first author with L. Zacchini.
3. Preliminaries
This section has two subsections, 3.1 on basic lemmas related to fractional integrals, and 3.2 on dyadic cubes.
3.1. Basic lemmas about fractional integrals
Here, we provide the proofs of the lemmas stated in Section 2, plus some additional ones to make those proofs more streamlined.
We begin with the sufficient condition for strong non-degeneracy stated in Lemma 2.5:
Proof of Lemma 2.5.
The non-empty annulus assumption implies that, for every point $x_{0}\in X$, radius $r>0$, and parameter $A>0$ yet to be chosen, there is a point $y_{0}\in B(x_{0},\overline{C}Ar)\setminus B(x,Ar)$. We will show that any such point satisfies properties (2.8) and (2.9) provided that $A>0$ is large enough (independently of $x_{0}$ and $r$).
Assumption (2.10) implies in particular that $K(x,y)>0$ for all $x\neq y$, and hence the bound concerning the argument in (2.8) and (2.9) is trivial with $v=1$ and $\eta=0$. As for the size bounds in (2.8) and (2.9), we note that both $\rho(x,y)\sim\rho(x_{0},y_{0})\approx Ar$ and $V(x,y)\sim V(x_{0},y_{0})\sim V(x_{0},Ar)$ for all $x,x_{0},y,y_{0}$ as in Definition 2.3 of strong non-degeneracy, as soon as $A$ is large enough. Fixing such an $A$, we then have $Ar\sim r$ and $V(x_{0},Ar)\approx V(x_{0},r)$, which implies that
$K(x,y)\sim\frac{\phi(x,y)}{V(x,y)}\sim\frac{\phi(x_{0},y_{0})}{V(x_{0},y_{0})}\sim\frac{\phi(x_{0},r)}{V(x_{0},r)},$
and the case of $K(y,x)$ is entirely analogous. ∎
To support the results about the well-definedness of $Tf$ and $[b,T]f$, we first give:
Lemma 3.1.
Let $\phi$ and KKăsatisfy (2.6) and (2.7). If $f\in L^{1}_{\operatorname{loc}}(X)$, then
(3.1) $\int_{B(x,r)}|K(x,y)f(y)|d\mu(y)\lesssim\phi(x,r)Mf(x),$
where $M$ is the Hardy–Littlewood maximal operator.
Proof..
In the metric case $\phi(x,y)=\rho(x,y)^{\alpha}$, we have
$\begin{split}\int_{B(x,r)}|K(x,y)f(y)|d\mu(y)&\sim\sum_{k=0}^{\infty}\int_{B(x,2^{-k}r)\setminus B(x,2^{-k-1}r)}\frac{(2^{-k}r)^{\alpha}}{V(x,2^{-k}r)}|f(y)|d\mu(y)\\ &\leq\sum_{k=0}^{\infty}(2^{-k}r)^{\alpha}Mf(x)\sim r^{\alpha}Mf(x).\end{split}$
In the volumic case $\phi(x,y)=V(x,y)^{\varepsilon}$, we choose a decreasing sequence $(r_{k})_{k=0}^{K}$, where $K\in\mathbb{N}\cup\{\infty\}$,ărecursively as follows: Let $r_{0}:=r$. Given $r_{k}$, we look for a $j\in\mathbb{N}$ such that $V(x,2^{-j}r_{k})<\frac{1}{2}V(x,r_{k})$. If no such $j\in\mathbb{N}$ exists, it means that $V(x,2^{-j}r_{k})\geq\frac{1}{2}V(x,r_{k})$ for all $j\in\mathbb{N}$, and hence $\mu(\{x\})=\lim_{k\to\infty}V(x,2^{-j}r_{k})\geq\frac{1}{2}V(x,r_{k})$. In this case, the process stops at this finite $k=:K$; note that this only happens if $\mu(\{x\})>0$. Otherwise, we let $r_{k+1}:=2^{-j}r_{k}$. Then, by construction,
$V(x,r_{k+1})<\frac{1}{2}V(x,r_{k})\leq V(x,2r_{k+1})\leq CV(x,r_{k+1})$
for allăk<Kk<K, and $\mu(\{x\})\leq V(x,r_{K})\leq 2\mu(\{x\})$. Thus we have the following, where the term involving $B(x,r_{K})$ is omitted if $K=\infty$:
$\begin{split}&\int_{B(x,r)}|K(x,y)f(y)|d\mu(y)\\ &=\sum_{k=0}^{K-1}\int_{B(x,r_{k})\setminus B(x,r_{k+1})}V(x,y)^{\varepsilon-1}d\mu(y)+\int_{B(x,r_{K})}V(x,y)^{\varepsilon-1}|f(y)|d\mu(y)\\ &\lesssim\sum_{k=0}^{K-1}V(x,r_{k+1})^{\varepsilon-1}\int_{B(x,r_{k})}|f(y)|d\mu(y)+\mu(\{x\})^{\varepsilon-1}\int_{B(x,r_{K})}|f(y)|d\mu(y)\\ &\sim\Big(\sum_{k=0}^{K-1}V(x,r_{k})^{\varepsilon}+V(x,r_{K})^{\varepsilon}\Big)Mf(x)\leq\sum_{k=0}^{K}(2^{-k}V(x,r))^{\varepsilon}Mf(x)\sim V(x,r)^{\varepsilon}Mf(x).\end{split}$
Thus, in both cases, we obtain (3.1). ∎
We can now provide the proof of Lemma 2.8 that we already stated in Section 2.
Proof of Lemma 2.8.
(i): Given $x\in X$, the assumption implies that $f$ is supported in $B(x,r)$ for some $r$. Hence the integral in (2.11) can be restricted to $B(x,r)$, and Lemma 3.1 guarantees that this integral exists provided that $Mf(x)<\infty$. By the weak (1,1)(1,1)ăinequality of $M$, this happens at almost every $x\in X$.
(ii): We can choose some $B=B(x_{0},r)$ such that both $f$ and $g$ are supported in $B$. Moreover, if $x,y\in B(x_{0},r)$, then $y\in B(x,2A_{0}r)$. It then follows from Lemma 3.1 that $|Tf(x)g(x)|\lesssim\phi(x,2A_{0}r)Mf(x)|g(x)|\lesssim\phi(x_{0},r)Mf(x)|g(x)|$. If $p>1$, then $Mf\in L^{p}(X)$, and hence this product is integrable. On the other hand, writing out the double integral defining $\langle Tf,g\rangle$, we also have the alternative bound
$\begin{split}\int_{X}|f(y)|\int_{B(y,2A_{0}r)}|K(x,y)||g(x)|d\mu(x)d\mu(y)&\lesssim\int_{X}|f(y)|\phi(y,2A_{0}r)Mg(y)d\mu(y)\\ &\lesssim\phi(x_{0},r)\int_{X}|f(y)|Mg(y)d\mu(y),\end{split}$
noting that the conditions (2.6) and (2.7) are, up to constants, symmetric in $x$ andăyy, so that Lemma 3.1 also applies with the roles of these variables interchanged. If $p^{\prime}>1$, then $Mg\in L^{p^{\prime}}(X)$, and the product above is integrable. Since at least one of $p,p^{\prime}$ is greater than 11, we can always apply at least one of these alternative bounds, both of which lead to the same result (2.12). ∎
3.2. Dyadic cubes
Some of our results will make use of the notion of systems of dyadic cubes in a space of homogeneous type $(X,\rho,\mu)$. Recall that the standard system of dyadic cubes on the Euclidean space $\mathbb{R}^{d}$ is defined as
$\mathscr{D}:=\{2^{-k}([0,1)^{d}+m):k\in\mathbb{Z},m\in\mathbb{Z}^{d}\}.$
The fundamental properties of these cubes are that any two of them are either disjoint or one is contained in the other, and that the cubes of a given size partition all space. As for general spaces of homogeneous type, a more general construction was first provided by Christ [8] and elaborated by Hytönen and Kairema [24], as follows.
Definition 3.2.
A system of dyadic cubes $\mathscr{D}$, on the space of homogeneous type $(X,\rho,\mu)$ is a collection
$\mathscr{D}=\bigcup_{k\in\mathbb{Z}}\mathscr{D}_{k},$
where
(1) for each $k\in\mathbb{Z}$, there is a disjoint union $X=\bigcup_{Q\in\mathscr{D}_{k}}Q$;
(2) each $\mathscr{D}_{k+1}$ refines the previous $\mathscr{D}_{k}$;
(3) for parameters $\delta\in(0,1)$ and $0<\widetilde{c_{0}}\leq\widetilde{C_{0}}$$\infty$, each $Q\in\mathscr{D}_{k}$ is essentially a ball of size ${\delta}^{k}$, in the sense that, for some “centre” $z_{Q}\in X$,
(3.2) $B(z_{Q},\widetilde{c_{0}}\delta^{k})\subseteq Q\subseteq B(z_{Q},\widetilde{C_{0}}\delta^{k}):=B_{Q}.$
We denote by $l(Q):=\delta^{k}$ the “side length” of $Q\in\mathscr{D}_{k}$.
Next, we will use the following notations from [23] to describe the relationship of different levels of dyadic cubes in Section 4. For $Q\in\mathscr{D}$, we denote by $Q^{[1]}$ the minimal $R\in\mathscr{D}$ such that $Q\subsetneq R$. We refer to $Q^{[1]}$ as the strict parent of $Q$ and $Q^{[1]}$ exists unless $Q=X$. Denote $Q^{[0]}:=Q$ and $Q^{[j]}:=(Q^{[j-1]})^{[1]}$ for each $j\in\mathbb{Z}$. By the equivalent size between one cube and its strict parent in Lemma 6.7, [23]: there exists constants $1<c\leq C<\infty$ such that for all $Q\in\mathscr{D}\setminus\{X\}$,
$c\mu(Q)\leq\mu(Q^{[1]})\leq C\mu(Q).$
By iterating the above estimate, it implies the following useful result: For every $\gamma>0$,
(3.3) $\sum_{k=0}^{\infty}\frac{1}{\mu(Q^{[k]})^{\gamma}}\leq\sum_{k=0}^{\infty}\frac{1}{c^{\gamma k}\mu(Q)^{\gamma}}=\frac{c^{\gamma}}{c^{\gamma}-1}\frac{1}{\mu(Q)^{\gamma}}.$
If the space of homogeneous type $(X,\rho,\mu)$ has a lower dimension $d>0$, then
$l(Q^{[k]})\sim\delta^{-k}l(Q),\,\,\,\mu(Q^{[k]})\gtrsim\delta^{-kd}\mu(Q),$
and hence for any $s,t>0$ satisfying $s<dt$,
(3.4) $\sum_{k=0}^{\infty}\frac{l(Q^{[k]})^{s}}{\mu(Q^{[k]})^{t}}\lesssim\frac{l(Q)^{s}}{\mu(Q)^{t}}\sum_{k=0}^{\infty}\delta^{k(dt-s)}\lesssim\frac{l(Q)^{s}}{\mu(Q)^{t}}.$
4. The upper bounds for the fractional commutators
For $p\in(2,\infty)$, Hytönen [23] obtained $S^{p}$ estimates for a class of general integral operators on spaces of homogeneous type, by extending an idea from Janson and Wolff [27] on the Euclidean space $\mathbb{R}^{d}$. For $p=2$ (Hilbert-Schmidt operators), it is classical. Based on these results, we give the following $S^{p}$ estimates for all $p\in[2,\infty)$. Let $(X,\rho,\mu)$ be a $\sigma$-finite measure space.
Proposition 4.1.
Let $(X,\rho,\mu)$ be a space of homogeneous type and $L$ be a kernel satisfying the size condition for any $x,y\in X$,
(4.1) $|L(x,y)|\lesssim\frac{1}{V(x,y)}.$
Then for any measurable function $B$ on $X\times X$ and exponent $p\in[2,\infty)$, the integral operator $\mathcal{I}_{BL}$:
$\mathcal{I}_{BL}f(x)=\int_{X}B(x,y)L(x,y)f(y)d\mu(y),\,\,\,\forall x\in X,$
satisfies
(4.2) $\|\mathcal{I}_{BL}\|_{S^{p}}\lesssim\|B\|_{L^{p}(V^{-2})}:=\left(\iint_{X\times X}\frac{|B(x,y)|^{p}}{V(x,y)^{2}}d\mu(x)d\mu(y)\right)^{\frac{1}{p}}.$
Proof..
For $p\in(2,\infty)$, this is [23, Proposition 5.8]. For $p=2$, an explicit formula of the Hilbert-Schmidt norm of integral operators (see e.g. [34, Theorem 2.11]) and (4.1) give that
$\|\mathcal{I}_{BL}\|_{S^{2}}=\|BL\|_{L^{2}}\lesssim\|BV^{-1}\|_{L^{2}}=\|B\|_{L^{2}(V^{-2})}.$
Thus, we complete the proof of (4.1). ∎
From Proposition 4.1, one immediately has the following two corollaries involving the Schatten class of the fractional commutators.
Corollary 4.2.
Let $p\in[2,\infty)$, $\varepsilon\in(0,1)$ and $\alpha\in(0,\infty)$. Let $(X,\rho,\mu)$ be a space of homogeneous type. Let $T\in\{T_{\varepsilon},\widetilde{T}_{\alpha}\}$ be a $\phi$-fractional integral with kernel $K$. Then the corresponding fractional commutator
$b,T]f(x)=\int_{X}[b(x)-b(y)]K(x,y)f(y)d\mu(y)$
belongs to the Schatten class $S^{p}$, provided that the respective right-hand side below is finite, and we have the estimates
(4.3) $\|[b,T_{\varepsilon}]\|_{S^{p}}\leq\|b\|_{\dot{B}_{p}^{\frac{1}{p}-\varepsilon}(\mu)}=\left(\int_{X}\int_{X}\frac{|b(x)-b(y)|^{p}}{V(x,y)^{1-p\varepsilon}}\cdot\frac{d\mu(x)d\mu(y)}{V(x,y)}\right)^{\frac{1}{p}},$
and
(4.4) $\|[b,\widetilde{T_{\alpha}}]\|_{S^{p}}\leq\|b\|_{\widetilde{B}_{p}^{\alpha}(\mu)}=\left(\int_{X}\int_{X}\frac{|b(x)-b(y)|^{p}\rho(x,y)^{p\alpha}}{V(x,y)}\cdot\frac{d\mu(x)d\mu(y)}{V(x,y)}\right)^{\frac{1}{p}}.$
Proof..
Taking $L(x,y)=\frac{1}{V(x,y)}$ and $B(x,y)=(b(x)-b(y))K(x,y)\cdot V(x,y)$ in (4.2), it follows from Proposition 4.1 that
$\left\|[b,T]\right\|_{S^{p}}=\|\mathcal{I}_{BL}\|_{S^{p}}\lesssim\left(\iint_{X\times X}|b(x)-b(y)|^{p}\frac{\phi(x,y)^{p}}{V(x,y)^{p}}V(x,y)^{p-2}d\mu(x)d\mu(y)\right)^{\frac{1}{p}},$
where
$\frac{\phi(x,y)^{p}}{V(x,y)^{p}}V(x,y)^{p-2}=\frac{\phi(x,y)^{p}}{V(x,y)^{2}}=\begin{cases}V(x,y)^{p\varepsilon-2},&\text{if }\phi(x,y)=V(x,y)^{\varepsilon},\\ \rho(x,y)^{p\alpha}V(x,y)^{-2},&\text{if }\phi(x,y)=\rho(x,y)^{\alpha}.\end{cases}$
Then we get (4.3) and (4.4), as desired. ∎
5. Characterization for fractional oscillation spaces and Besov spaces
If $p\in(1,\infty)$, Hytönen [23] described Besov spaces $\dot{B}_{p}(\mu)$ by oscillation spaces $\mathrm{Osc}^{p}(\mu)$ (see [23], Proposition 1.29), namely
$\|b\|_{\dot{B}_{p}(\mu)}:=\left(\int_{X}\int_{X}\frac{|b(x)-b(y)|^{p}}{V(x,y)^{2}}d\mu(x)d\mu(y)\right)^{\frac{1}{p}}$
$\thickapprox\|b\|_{\mathrm{Osc}^{p}(\mu)}:=\Big\|\Big\{\fint_{B_{Q}}|b-\langle b\rangle_{B_{Q}}|d\mu\Big\}_{Q\in\mathscr{D}}\Big\|_{\ell^{p}}.$
For the fractional version, an analogous is stated below in Propositions 5.2. First, we give a useful characterisation from [23, Lemma 11.4]:
Lemma 5.1 ([23], Lemma 11.4).
For any $p\in(0,\infty)$,
(5.1) $\inf_{c}\Big(\fint_{B}|b-c|^{p}d\mu\Big)^{\frac{1}{p}}$ $\sim\Big(\fint_{B}\fint_{B}|b(x)-b(y)|^{p}d\mu(x)d\mu(y)\Big)^{\frac{1}{p}}$
$\sim\Big(\fint_{B}|b(x)-\langle b\rangle_{B}|^{p}d\mu(x)\Big)^{\frac{1}{p}},\,\,\text{if}\,\,p\in[1,\infty).$
Consistently with (2.6), for any cube $Q$, we denote
(5.2) $\phi(Q)\in\{\mu(B_{Q})^{\varepsilon},l(Q)^{\alpha}\}.$
For $p\in(1,\infty)$, we denote
(5.3) $\|b\|_{{\mathrm{Osc}}_{\phi}^{p}(\mu)}:=\Big\|\Big\{{\phi(Q)}\fint_{B_{Q}}|b-\langle b\rangle_{B_{Q}}|d\mu\Big\}_{Q\in\mathscr{D}}\Big\|_{\ell^{p}}.$
Proposition 5.2.
Let $(X,\rho,\mu)$ be a space of homogeneous type with a lower dimension $d>0$ and $b\in L_{\mathrm{loc}}^{1}(\mu)$. Let $\mathscr{D}$ be a system of dyadic cubes on $(X,\rho,\mu)$ in the sense of Definition 3.2 and, for each $Q\in\mathscr{D}$, let $B_{Q}$ be a ball centred at $Q$ and of radius $c\cdot l(Q)$ for a constant $c$ that only depends on the space $X$. Let $p\in(1,\infty)$, and $\phi$ be as in (2.6) and (5.2) with $\varepsilon\in(0,1/p)$ and $\alpha\in(0,d/p)$. Then
$\left\|b\right\|_{\mathbb{B}_{p}(\phi,\mu)}\sim\|b\|_{{\mathrm{Osc}}_{\phi}^{p}(\mu)}.$
Proposition 5.2 will be a consequence of the following Propositions 5.3 and 5.4. We note that the range of the admissible parameters is larger in Proposition 5.3.
Proposition 5.3.
Suppose that the space of homogeneous type $(X,\rho,\mu)$ has a lower dimension $d\in(0,\infty)$. Let $p\in(1,\infty)$ and let $\phi$ăbe as in (2.6) and (5.2) with $\varepsilon\in(0,\frac{2}{p})$ and $\alpha\in(0,\frac{2d}{p})$. For all measurable $b$ on $X$, we have
$\left\|b\right\|_{\mathbb{B}_{p}(\phi,\mu)}\sim\Big\|\Big\{\phi(Q)^{\varepsilon}\inf_{c}\Big(\fint_{B_{Q}}|b-c|^{p}d\mu\Big)^{\frac{1}{p}}\Big\}_{Q\in\mathscr{D}}\Big\|_{\ell^{p}}.$
Proof..
By (5.1), it follows that
(5.4) $\Big\|\Big\{\phi(Q)\inf_{c}\Big(\fint_{B_{Q}}|b-c|^{p}d\mu\Big)^{\frac{1}{p}}\Big\}_{Q\in\mathscr{D}}\Big\|_{\ell^{p}}^{p}$
$\sim\Big\|\Big\{\phi(Q)\Big(\fint_{B_{Q}}\fint_{B_{Q}}|b(x)-b(y)|^{p}d\mu(x)d\mu(y)\Big)^{\frac{1}{p}}\Big\}_{Q\in\mathscr{D}}\Big\|_{\ell^{p}}^{p}$
$=\sum_{Q\in\mathscr{D}}{\phi(Q)^{p}}\fint_{B_{Q}}\fint_{B_{Q}}|b(x)-b(y)|^{p}d\mu(x)d\mu(y)$
$=\int_{X}\int_{X}|b(x)-b(y)|^{p}\sum_{Q\in\mathscr{D}}1_{B_{Q}}(x)1_{B_{Q}}(y)\cdot\frac{\phi(Q)^{p}}{{\mu(B_{Q})}^{2}}d\mu(x)d\mu(y).$
Next, we consider the sum in the above formula. On the one hand, there is a dyadic cube $Q$ which contains $x$ and $\ell(Q)\sim\rho(x,y)$. Under this condition, we obtain that the corresponding $B_{Q}$ contains both $x$ and $y$, satisfying $\mu(B_{Q})\sim V(x,y)$. Hence, for this $Q$,
$\phi(Q)=\left.\begin{cases}\mu(B_{Q})^{\varepsilon}&\sim\ V(x,y)^{\varepsilon}\\ l(Q)^{\alpha}&\sim\ \rho(x,y)^{\alpha}\end{cases}\right\}=\phi(x,y),\qquad\frac{\phi(Q)^{p}}{\mu(B_{Q})^{2}}\sim\frac{\phi(x,y)^{p}}{V(x,y)^{2}}$
according to the two cases in (2.6) and (5.2), and thus
(5.5) $\sum_{Q\in\mathscr{D}}1_{B_{Q}}(x)1_{B_{Q}}(y)\cdot\frac{\phi(Q)^{p}}{{\mu(B_{Q})}^{2}}\gtrsim\frac{\phi(x,y)^{p}}{V(x,y)^{2}}.$
On the other hand, for any $x,y\in B_{Q}$ with $x\neq y$, it follows that $\ell(Q)\gtrsim\rho(x,y)$. For this type of dyadic cube $Q$, it contains a minimal cube belonging to the same type and there are at most boundedly many such minimal cubes thanks to the doubling property. Then every $Q$ appearing in the sum is a dyadic ancestor $P^{[k]}$ of some minimal cube $P$. Now we consider the following two cases.
Case 1: $\phi(Q)=\mu(B_{Q})^{\varepsilon}$ with $\varepsilon\in(0,\frac{2}{p})$. Taking $Q=P$ and $\gamma=2-p\varepsilon>0$ in (3.3), we get
$\sum_{k=0}^{\infty}\frac{1}{\mu(P^{[k]})^{2-p\varepsilon}}\lesssim\frac{1}{\mu(P)^{2-p\varepsilon}}\sim\frac{1}{V(x,y)^{2-p\varepsilon}}.$
Summing over the boundedly many minimal $P$, we conclude that
(5.6) $\sum_{Q\in\mathscr{D}}\frac{1_{B_{Q}}(x)1_{B_{Q}}(y)}{{\mu(B_{Q})}^{2-p\varepsilon}}\lesssim\frac{1}{V(x,y)^{2-p\varepsilon}}=\frac{\phi(x,y)^{p}}{V(x,y)^{2}}.$
Case 2: $\phi(Q)=l(Q)^{\alpha}$ with $\alpha\in(0,\frac{2d}{p})$. Taking $Q=P$, $s=p\alpha$ and $t=2$ in (3.4), noting that $p\alpha<p\frac{2d}{p}=2d$, we get
$\sum_{k=0}^{\infty}\frac{l(P^{[k]})^{p\alpha}}{\mu(P^{[k]})^{2}}\lesssim\frac{l(P)^{p\alpha}}{\mu(P)^{2}}\thickapprox\frac{\rho(x,y)^{p\alpha}}{V(x,y)^{2}}.$
Summing over the boundedly many minimal $P$, we conclude that for any $\alpha\in(0,\frac{2d}{p})$,
(5.7) $\sum_{Q\in\mathscr{D}}1_{B_{Q}}(x)1_{B_{Q}}(y)\cdot\frac{{l(Q)}^{p\alpha}}{{\mu(B_{Q})}^{2}}\lesssim\frac{\rho(x,y)^{p\alpha}}{V(x,y)^{2}}=\frac{\phi(x,y)^{p}}{V(x,y)^{2}}$
By (5.5) through (5.7), we find that
$\sum_{Q\in\mathscr{D}}1_{B_{Q}}(x)1_{B_{Q}}(y)\cdot\frac{\phi(Q)^{p}}{{\mu(B_{Q})}^{2}}\sim\frac{\phi(x,y)^{p}}{V(x,y)}.$
Substituting this into (5.4), we finish the proof of Proposition 5.3. ∎
Proposition 5.4.
Suppose that the space of homogeneous type $(X,\rho,\mu)$ has a lower dimension $d\in(0,\infty)$. Let $p\in(1,\infty)$ and $\phi$ be as in (2.6) and (5.2) with $\varepsilon\in(0,\frac{1}{p})$ and $\alpha\in(0,\frac{d}{p})$. For the oscillatory norm $\left\|\cdot\right\|_{\mathrm{Osc}_{\phi}^{p}(\mu)}$ defined in (5.3) and any $b\in L_{\mathrm{loc}}^{1}(\mu)$, we have
(5.8) $\Big\|\Big\{\phi(Q)\inf_{c}\Big(\fint_{B_{Q}}|b-c|^{p}d\mu\Big)^{\frac{1}{p}}\Big\}_{Q\in\mathscr{D}}\Big\|_{\ell^{p}}\sim\|b\|_{\mathrm{Osc}_{\phi}^{p}(\mu)}.$
Proof..
For a given $Q\in\mathscr{D}$ and any $r\in(0,\infty)$ and $\varepsilon\in(0,1)$, we define
$\lambda_{r}^{\phi}(Q)$ $:=\inf_{c}\phi(Q)\Big(\fint_{Q}|b-c|^{r}d\mu\Big)^{\frac{1}{r}}$
$=\phi(Q)\inf_{c}\Big(\fint_{Q}|b-c|^{r}d\mu\Big)^{\frac{1}{r}}=:\phi(Q)\lambda_{r}(Q)$
[…]
$\mathrm{Car}\lambda(P):=\frac{1}{\mu(P)}\sum_{Q\subseteq P}|\lambda_{Q}|\mu(Q).$
Taking $r=p$ and $s=1$ in [23, Proposition 11.3], we get for any given $Q\in\mathscr{D}$,
(5.9) $\lambda_{p}(Q)\lesssim\big(\mathrm{Car}(\lambda_{1})^{p}\big)^{\frac{1}{p}}(Q)=\Big(\frac{1}{\mu(Q)}\sum_{S\subseteq Q}\lambda_{1}(S)^{p}\mu(S)\Big)^{\frac{1}{p}}.$
It implies that
(5.10) $\|\{\lambda_{p}^{\phi}(Q)\}$ ${}_{Q\in\mathscr{D}}\|_{\ell^{p}}^{p}\lesssim\|\{\phi(Q)\big(\mathrm{Car}(\lambda_{1})^{p}\big)^{\frac{1}{p}}(Q)\}_{Q\in\mathscr{D}}\|_{\ell^{p}}^{p}=\sum_{Q\in\mathscr{D}}\phi(Q)^{p}\mathrm{Car}(\lambda_{1})^{p}(Q)$
$=\sum_{Q\in\mathscr{D}}\phi(Q)^{p}\sum_{S\subseteq Q}\frac{\mu(S)}{\mu(Q)}\lambda_{1}(S)^{p}=\sum_{S\in\mathscr{D}}\lambda_{1}(S)^{p}\sum_{Q\supseteq S}\frac{\mu(S)}{\mu(Q)}\cdot\phi(Q)^{p}$
$=\sum_{S\in\mathscr{D}}\lambda_{1}(S)^{p}\sum_{k=0}^{\infty}\frac{\mu(S)}{\mu(S^{[k]})}\cdot\phi(S^{[k]})^{p},$
where $S^{[k]}$ is the $k$-th strict dyadic ancestor. Similarly as in the proof of Proposition 5.3, we consider the two cases.
Case 1: $\phi(Q)=\mu(B_{Q})^{\varepsilon}$ with $\varepsilon\in(0,\frac{1}{p})$. Taking $Q=S$ and $\gamma=1-p\varepsilon>0$ in (3.3), we get
(5.11) $\sum_{k=0}^{\infty}\frac{\mu(S)}{\mu(S^{[k]})}\cdot\phi(S^{[k]})^{p}=\sum_{k=0}^{\infty}\frac{\mu(S)}{\mu(S^{[k]})^{1-p\varepsilon}}\lesssim\frac{\mu(S)}{\mu(S)^{1-p\varepsilon}}=\mu(S)^{p\varepsilon}\sim\phi(S)^{p}.$
Case 2: $\phi(Q)=l(Q)^{\alpha}$ with $\alpha\in(0,\frac{d}{p})$. Taking $Q=S$, $s=p\alpha$ and $t=1$ in (3.4), noting that $p\alpha<p\frac{d}{p}=d\cdot 1$ as required to apply (3.4), it follows that
(5.12) $\sum_{k=0}^{\infty}\frac{\mu(S)}{\mu(S^{[k]})}\cdot\phi(S^{[k]})^{p}=\sum_{k=0}^{\infty}\frac{\mu(S)}{\mu(S^{[k]})}\cdot l(S^{[k]})^{p\alpha}\lesssim\mu(S)\frac{l(S)^{p\alpha}}{\mu(S)}=l(S)^{p\alpha}=\phi(S)^{p}.$
Thus, (5.11) and (5.12) yield that
$\sum_{k=0}^{\infty}\frac{\mu(S)}{\mu(S^{[k]})}\cdot\phi(S^{[k]})^{p}\lesssim\phi(S)^{p}.$
Substituting this into (5.10), we get
$\|\{\lambda_{p}^{\phi}(Q)\}_{Q\in\mathscr{D}}\|_{\ell^{p}}^{p}\lesssim\sum_{S\in\mathscr{D}}\lambda_{1}(S)^{p}\phi(S)^{p}=\|\{\lambda_{1}^{\phi}(Q)\}_{Q\in\mathscr{D}}\|_{\ell^{p}}^{p}\sim\|b\|_{\mathrm{Osc}_{\phi}^{p}(\mu)},$
where we used (5.1) in the last equality $\sim$”.
This proves $\lesssim$” in (5.8), while $\gtrsim$” is clear. ∎
6. The complex median method of Wei–Zhang revisited
Two methods for proving lower bounds for commutators in the recent literature are the “approximate weak factorisation” from [22], and versions of the so-called median method. Until recently, this second approach was restricted to real-valued functions, for which the median is conventionally defined. However, Wei and Zhang [36] recently demonstrated that, with suitable modifications, the median method can also be extended to work with complex-valued functions. Once their [36, Theorem 1.5] on the existence of the complex median is proved, it can be conveniently applied as a black box; however, the proof of [36, Theorem 1.5] itself is not so easy. For this reason, we present below a variant of this theorem, with a much simpler proof and roughly the same scope of applicability.
For $u\in\mathbb{C}$ with $|u|=1$ and $\theta\in[0,2\pi]$, we denote by
$\begin{split}\Gamma(u,\theta)&:=\{z\in\mathbb{C}:\operatorname{Re\,}(\bar{u}z)\geq\cos(\tfrac{1}{2}\theta)|z|\},\\ \Gamma^{\circ}(u,\theta)&:=\{z\in\mathbb{C}:\operatorname{Re\,}(\bar{u}z)>\cos(\tfrac{1}{2}\theta)|z|\}\end{split}$
the closed and open cones in direction $u$ of total angle $\theta$ (thus $\frac{1}{2}\theta$ on either side of $u$).
Proposition 6.1.
Let $(N,\varepsilon)=(3,\frac{1}{4})$. Given a Borel probability $\nu$ on $\mathbb{C}$, there exist $m\in\mathbb{C}$ and $N$ closed cones $\{\Gamma_{j}\}_{j=1}^{N}$ of angle $\frac{2\pi}{N}$ each, together covering all $\mathbb{C}$, such that
$\nu(m+\Gamma_{j})\geq\varepsilon$
for every $j\in\{1,\ldots,N\}$.
Remark 6.2.
Recall that the median of a probability $\nu$ on $\mathbb{R}$ is a value $m\in\mathbb{R}$ such that both $\nu((-\infty,m])\geq\frac{1}{2}$ and $\nu([m,\infty))\geq\frac{1}{2}$, i.e., the mass of $\nu$ essentially equally divided on two sides of $m$, to the extent allowed by a possible point mass at $m$.
The value $m\in\mathbb{C}$ in Proposition 6.1 is a “quasi-median” of $\nu$ in the sense that the measure of $\nu$ is divided among each of the $N$ conical regions around $m$, with a fair (if not quite equal) share to each.
Proposition 6.1 also holds in each of the cases $(N,\varepsilon)\in\{(1,1),(2,\tfrac{1}{2}),(4,\tfrac{1}{16})\}$, where case $(N,\varepsilon)=(1,1)$ is the triviality $\nu(\mathbb{C})\geq 1$, while case $N=2$ and $\varepsilon=\frac{1}{2}$ is a simple extension of the existence of the median for the probability $\nu_{1}(A):=\nu(A\times\mathbb{R})$ on $\mathbb{R}$: if $m$ is a median of $\nu_{1}$, the required property is satisfied with the cones $\Gamma(\pm 1,\pi)$.
The nontrivial case $(N,\varepsilon)=(4,\frac{1}{16})$ is a recent result of Wei and Zhang, [36, Theorem 1.5]. Its proof is “elementary” but rather tedious, taking about 10 pages in [36, Section 6].
We will show that case $(N,\varepsilon)=(3,\frac{1}{4})$ holds with a much simpler proof, while the result has roughly the same scope of applicability as [36, Theorem 1.5] for commutator estimates, as we will see further below.
The proof below shows that the cones can be chosen to be in the standard orientation $\Gamma_{j}=\Gamma(e^{ij\frac{2\pi}{3}},\frac{2\pi}{3})$ (or, by a rotation, in any other prescribed orientation that we like). This is in contrast to [36, Theorem 1.5], where the orientation of the cones may depend on $\nu$.
Proof of Proposition 6.1.
We will identify $\mathbb{C}\simeq\mathbb{R}^{2}$ without explicit mention whenever convenient. Let $t$ be a median of $\nu_{2}:E\subset\mathbb{R}\mapsto\nu(\mathbb{R}\times E)$. Hence both
$\nu(z:\operatorname{Im\,}z\geq t)\geq\frac{1}{2},\qquad\nu(z:\operatorname{Im\,}z\leq t)\geq\frac{1}{2}.$
Let
$\sigma_{0}:=\sup\{\sigma:\nu(\sigma+it+\Gamma(1,\tfrac{2\pi}{3}))\geq\tfrac{1}{4}\}.$
We claim that $m:=\sigma_{0}+it$ satisfies the required property with the cones $\Gamma_{0}:=\Gamma(1,\frac{2\pi}{3})$ and $\Gamma_{\pm}:=\Gamma(u_{\pm},\frac{2\pi}{3})$, where $u_{\pm}:=e^{\pm i\frac{2\pi}{3}}$. Since
$m+\Gamma_{0}=\bigcap_{\sigma<\sigma_{0}}(\sigma+it+\Gamma_{0}),\qquad m+\Gamma_{0}^{\circ}=\bigcup_{\sigma>\sigma_{0}}(\sigma+it+\Gamma_{0}),$
it follows from the continuity of measure and the definition of supremum that
$\nu(m+\Gamma_{0})\geq\tfrac{1}{4},\qquad\nu(m+\Gamma^{\circ}_{0})\leq\tfrac{1}{4}.$
Noting that
$\mathbb{C}\setminus\Gamma_{0}^{\circ}\subset\Gamma_{+}\cup\{z:\operatorname{Im\,}z<0\},$
it follows that
$\tfrac{3}{4}\leq\nu(m+\mathbb{C}\setminus\Gamma^{\circ}_{0})\leq\nu(m+\Gamma_{+})+\nu(\{z:\operatorname{Im\,}z<t\})\leq\nu(m+\Gamma_{+})+\tfrac{1}{2}$
recalling that $t$ is a median of $\nu_{2}$. Hence $\nu(m+\Gamma_{+})\geq\tfrac{3}{4}-\tfrac{1}{2}=\tfrac{1}{4}$, and the proof of the similar estimate with $\Gamma_{-}$ in place of $\Gamma_{+}$ is entirely analogous. This completes the proof. ∎
The following result implements the “median method” in the present setting: With the help of the “median” from Proposition 6.1, we dominate the oscillations of $b$ by the action of a commutator $[b,T]$ on suitable test functions.
Proposition 6.3.
Let $K:X\times X\to\mathbb{C}$ be a kernel, and let $B$ and $\tilde{B}$ be balls such that
(6.1) $\left|K(x,y)\right|\gtrsim\frac{\phi(B)}{\mu(B)},\qquad\left|\arg(\bar{v}K(x,y))\right|\leq\eta\leq\frac{\pi}{9}$
for some $v\in\mathbb{C}$ with $\left|v\right|=1$ and for all $x\in B$ and $y\in\tilde{B}$. If $b\in L^{1}_{\operatorname{loc}}(X)$, then there are measurable subsets $E\subset B$ and $F\subset\tilde{B}$ such that
$\inf_{c}\phi(B)\fint_{B}\left|b(x)-c\right|d\mu(x)\lesssim\frac{1}{\mu(\tilde{B})}\left|\langle 1_{E},[b,T]1_{F}\rangle\right|,$
where $T$ is the integral operator with kernel $K$.
Proof..
[…]
Since $\bigcup_{j=1}^{N}\Gamma_{j}=\mathbb{C}$, it follows that
(6.2) $\phi(B)\fint_{B}\left|b(x)-m\right|d\mu(x)\leq\sum_{j=1}^{N}\int_{B\cap\{b\in m+\Gamma_{j}\}}\frac{\phi(B)}{\mu(B)}\left|b(x)-m\right|d\mu(x).$
Now consider a fixed $\Gamma_{j}=\Gamma(u_{j},\frac{2\pi}{N})$. We claim that, for every $x\in B\cap\{b\in m+\Gamma_{j}\}$ and $y\in\tilde{B}\cap\{b\in m-\Gamma_{j}\}$, we have
(6.3) $\frac{\phi(B)}{\mu(B)}\left|b(x)-m\right|\lesssim\operatorname{Re\,}[\overline{u_{j}v}(b(x)-b(y))K(x,y)],$
where $v\in\mathbb{C}$ comes from the assumption (6.1)
To justify (6.3), let $b(x)=m+u_{j}t_{1}e^{i\phi_{1}}$ and $b(y)=m-u_{j}t_{2}e^{i\phi_{2}}$, where $t_{1}=\left|b(x)-m\right|$, $t_{2}=\left|b(y)-m\right|$, and $\left|\phi_{j}\right|\leq\theta=\frac{\pi}{N}$. By (6.1), we can also write $K(x,y)=vt_{0}e^{i\phi_{0}}$, where $t_{0}=\left|K(x,y)\right|\gtrsim\phi(B)\mu(B)^{-1}$. Hence
$\begin{split}\operatorname{Re\,}[\overline{u_{j}v}(b(x)-b(y))K(x,y)]&=\operatorname{Re\,}[\bar{u}_{j}((m+u_{j}t_{1}e^{i\phi_{1}})-(m-u_{j}t_{2}e^{i\phi_{2}}))\bar{v}vt_{0}e^{i\phi_{0}}]\\ &=\operatorname{Re\,}(t_{1}t_{0}e^{i(\phi_{1}+\phi_{0})}+t_{2}t_{0}e^{i(\phi_{2}+\phi_{0})})\\ &=t_{1}t_{0}\cos(\phi_{1}+\phi_{0})+t_{2}t_{0}\cos(\phi_{2}+\phi_{0}),\end{split}$
where
$\left|\phi_{j}+\phi_{0}\right|\leq\theta+\eta=\frac{\pi}{N}+\frac{\pi}{9}\leq\frac{\pi}{3}+\frac{\pi}{9}=\frac{4}{9}\pi<\frac{\pi}{2}.$
Since $\cos\phi=\cos\left|\phi\right|$ is a decreasing function of $\left|\phi\right|\in[0,\frac{\pi}{2}]$, it follows that
$\begin{split}\operatorname{Re\,}[\overline{u_{j}v}(b(x)-b(y))K(x,y)]&\geq(t_{1}+t_{2})t_{0}\cos(\tfrac{4}{9}\pi)\\ &\gtrsim t_{1}t_{0}\gtrsim\left|b(x)-m\right|\frac{\phi(B)}{\mu(B)}.\end{split}$
This proves (6.3).
We can now take the average of (6.3) over $y\in\tilde{B}\cap\{b\in m-\Gamma_{j}\}$. This gives
$\begin{split}\frac{\phi(B)}{\mu(B)}\left|b(x)-m\right|&\lesssim\fint_{\tilde{B}\cap\{b\in m-\Gamma_{j}\}}\operatorname{Re\,}[\overline{u_{j}v}(b(x)-b(y))K(x,y)]d\mu(y)\\ &\lesssim\operatorname{Re\,}\Big[\overline{u_{j}v}\fint_{\tilde{B}}(b(x)-b(y))K(x,y)1_{\tilde{B}\cap\{b\in m-\Gamma_{j}\}}(y)d\mu(y)\Big],\end{split}$
[…]
$\begin{split}&\phi(B)\fint_{B}\left|b(x)-m\right|d\mu(x)\\ &\lesssim\sum_{j=1}^{N}\operatorname{Re\,}\Big[\overline{u_{j}v}\int_{B}1_{B\cap\{b\in m+\Gamma_{j}\}}(x)\fint_{\tilde{B}}(b(x)-b(y))K(x,y)1_{\tilde{B}\cap\{b\in m-\Gamma_{j}\}}(y)d\mu(y)d\mu(x)\Big]\\ &=\mu(\tilde{B})^{-1}\sum_{j=1}^{N}\operatorname{Re\,}\Big[\overline{u_{j}v}\left\langle 1_{B\cap\{b\in m+\Gamma_{j}\}},[b,T]1_{\tilde{B}\cap\{b\in m-\Gamma_{j}\}}\right\rangle\Big].\end{split}$
The proof is completed by estimating the real part by the absolute value, and taking for $(E,F)$ the $(B\cap\{b\in m+\Gamma_{j}\},\tilde{B}\cap\{b\in m-\Gamma_{j}\})$ for which the largest value is achieved. ∎
Corollary 6.4.
Let $T$ be a strongly non-degenerate $\phi$-fractional integral. Then for all balls $B\subset X$, there exist subsets $E,F\subset B^{*}$, where $B^{*}=C\cdot B$ with a fixed dilation factor $C$, such that
$\phi(B)\int_{B}\left|b-\langle b\rangle_{B}\right|d\mu\lesssim\left|\langle 1_{E},[b,T]1_{F}\rangle\right|.$
Proof..
Consider a fixed ball $B=B(x_{0},r)$. By assumption (recall Definition 2.3), for some $y_{0}\in B(x_{0},\overline{C}Ar)\setminus B(x_{0},Ar)$, the kernel $K$ of $T$ satisfies either (2.8) or (2.9). Let us first assume (2.8) and denote $\tilde{B}:=B(y_{0},r)$. If $y\in\tilde{B}$, then
$\rho(y,x_{0})\leq A_{0}\rho(y,y_{0})+A_{0}\rho(y_{0},x_{0})<A_{0}(1+\overline{C}A)r,$
thus $\tilde{B}\subset B^{*}:=B(x_{0},A_{0}(1+\overline{C}A)r)=C\cdot B$ with $C=A_{0}(1+\overline{C}A)$, and symmetrically $B\subset C\cdot\tilde{B}$.
Note that (2.8) coincides with assumption (6.1) of Proposition 6.3; hence the conclusion of Proposition 6.3 gives us $E\subset B\subset B^{*}$ and $F\subset\tilde{B}\subset B^{*}$ such that
$\phi(B)\int_{B}\left|b-\langle b\rangle_{B}\right|d\mu\lesssim\inf_{c}\phi(B)\int_{B}\left|b-c\right|d\mu\lesssim\frac{\mu(B)}{\mu(\tilde{B})}\left|\langle 1_{E},[b,T]1_{F}\rangle\right|.$
Since $B\subset C\cdot\tilde{B}$, it follows from doubling that
[…]
and we get the claimed estimate in this first case.
Suppose then that (2.9) holds instead. Let $K^{*}(x,y)=K(y,x)$, and let $T^{*}g(u):=\int_{X}K^{*}(u,v)f(v)d\mu(v)$. Then $T^{*}$ is also a $\phi$-fractional integral operator, whose kernel satisfies the assumptions of the previous case in the ball $B$. Thus, by the previous case, we find sets $E,F\subset B^{*}$ such that
$\phi(B)\int_{B}\left|b-\langle b\rangle_{B}\right|\lesssim\left|\langle 1_{E},[b,T^{*}]1_{F}\rangle\right|.$
But
$\langle 1_{E},[b,T^{*}]1_{F}\rangle=\langle 1_{E},bT^{*}1_{F}-T^{*}(b1_{F})\rangle=\langle T(b1_{E}),1_{F}\rangle-\langle bT1_{E},1_{F}\rangle=-\langle 1_{F},[b,T]1_{E}\rangle,$
and hence we get the claimed estimate, only with the names of $E$ and $F$ interchanged. This completes the proof. ∎
7. The lower bounds for the fractional commutators
The main results of this section are lower bounds for the Schatten norm of fractional commutators $[b,T]$, by means of fractional oscillation norms that generalize those introduced in Section 5. Although our main concern in this paper is Schatten $S^{p}$ norms, we state and prove the following estimate for the more general Schatten–Lorentz $S^{p,q}$ănorms: on the one hand, this added generality comes essentially for free by the same argument; on the other hand, it may be useful elsewhere. We denote by $\ell^{p,q}$ the usual Lorentz sequence space (see e.g. [2]) and by $S^{p,q}=S^{p,q}(L^{2}(\mu))$ the space of compact operators on $L^{2}(\mu)$ with
$\left\|R\right\|_{S^{p,q}}:=\left\|\{a_{n}(R)\}_{n=0}^{\infty}\right\|_{\ell^{p,q}},$
where $a_{n}(R)$ are the approximation numbers (or singular values) as in (1.2).
Proposition 7.1.
Let $(X,\rho,\mu)$ăbe a space of homogeneous type with a system of dyadic cubes $\mathscr{D}$ in the sense of Definition 3.2 and, for each $Q\in\mathscr{D}$, let $B_{Q}$ be a ball centred at $Q$ and of radius $c\cdot l(Q)$ for a constant $c$ only depending on the space $X$. Let $q\in[1,\infty]$, $p\in(1,\infty)$, and $\phi$ be as in (2.6) and (5.2) with $\varepsilon\in(0,1)$ and $\alpha\in(0,\infty)$. Let $T$ be a strongly non-degenerate $\phi$-fractional integral operator. Then for all $b\in L_{\mathrm{loc}}^{1}(\mu)$, we have
(7.1) $\|b\|_{\mathrm{Osc}_{\phi}^{p,q}(\mu)}:=\Big\|\Big\{\phi(Q)\fint_{B_{Q}}|b-\langle b\rangle_{B_{Q}}|d\mu\Big\}_{Q\in\mathscr{D}}\Big\|_{\ell^{p,q}}.\lesssim\|[b,T]\|_{S^{p,q}},$
In particular, when $p=q\in(1,\infty)$, we have
(7.2) $\|b\|_{\mathrm{Osc}_{\phi}^{p}(\mu)}=\|b\|_{\mathrm{Osc}_{\phi}^{p,p}(\mu)}\lesssim\|[b,T]\|_{S^{p,p}}=\|[b,T]\|_{S^{p}}.$
To prove Proposition 7.1, motivated by the ideas of [23] which go back to [33], we consider the bi-sublinear maximal operator $\mathcal{M}$ for a sequence $\{(e_{Q},h_{Q})\}_{Q\in\mathscr{D}}$ of pairs of functions,
(7.3) $\mathcal{M}:(f,g)\mapsto\sup_{Q\in\mathscr{D}}1_{Q}\frac{|\langle f,e_{Q}\rangle\langle g,h_{Q}\rangle|}{\mu(Q)},$
and two related maximal operators
$\mathcal{M}^{1}f:=\sup_{Q\in\mathscr{D}}1_{Q}\frac{|\langle f,e_{Q}\rangle|}{\mu(Q)^{\frac{1}{2}}},\quad\mathcal{M}^{2}g:=\sup_{Q\in\mathscr{D}}1_{Q}\frac{|\langle g,h_{Q}\rangle|}{\mu(Q)^{\frac{1}{2}}}.$
Obviously, $\mathcal{M}(f,g)\leq\mathcal{M}^{1}f\cdot\mathcal{M}^{2}g$. If the sequence of functions $\{(e_{Q},h_{Q})\}_{Q\in\mathscr{D}}$ satisfies
(7.4) $|e_{Q}|+|h_{Q}|\lesssim\mu(Q)^{-\frac{1}{2}}1_{B_{Q}^{*}},$
where $B_{Q}^{*}=c\cdot B_{Q}$ is a concentric extension of $B_{Q}$ for any fixed constant $c$, then
$\mathcal{M}^{1}f\lesssim Mf,\qquad\mathcal{M}^{2}g\lesssim Mg,$
where $M$ is the Hardy-Littlewood maximal operator. Under the assumption (7.4), by Hölder’s inequality and the $(L^{2}(\mu),L^{2}(\mu))$ boundedness of the maximal operator $M$, we obtain
$\|\mathcal{M}(f,g)\|_{L^{1}(\mu)}\leq\|\mathcal{M}^{1}f\|_{L^{2}(\mu)}\|\mathcal{M}^{2}g\|_{L^{2}(\mu)}\lesssim\|f\|_{L^{2}(\mu)}\|g\|_{L^{2}(\mu)}.$
Thus, the bi-sublinear maximal operator $\mathcal{M}$ for a sequence of functions $\{(e_{Q},h_{Q})\}_{Q\in\mathscr{D}}$ satisfying (7.4), is bounded from $L^{2}(\mu)\times L^{2}(\mu)$ to $L^{1}(\mu)$.
We will also need the following result, which is [23, Corollary 7.7]:
Proposition 7.2.
Let $p\in(1,\infty)$ and $q\in[1,\infty]$. For all sequences $\{(e_{Q},h_{Q})\}_{Q\in\mathscr{D}}$ and their related maximal operator (7.3), we have the following estimate for operators $A\in S^{p,q}(L^{2}(\mu))$:
$\|\{\langle Ae_{Q},h_{Q}\rangle\}_{Q\in\mathscr{D}}\|_{l^{p,q}(\mathscr{D})}\lesssim\|\mathcal{M}\|_{L^{2}(\mu)\times L^{2}(\mu)\rightarrow L^{1}(\mu)}\|A\|_{S^{p,q}(L^{2}(\mu))}.$
We are now prepared to give:
Proof of Proposition 7.1.
For each $Q\in\mathscr{D}$, we apply Corollary 6.4 with $B=B_{Q}$. This provides us with subsets $E_{Q},F_{Q}\subset B_{Q}^{*}$ such that
$m_{b}^{\phi}(B_{Q}):=\phi(Q)\fint_{B_{Q}}|b-\langle b\rangle_{B_{Q}}|d\mu\lesssim|\langle[b,T](\frac{1_{E_{Q}}}{\mu(B_{Q})^{\frac{1}{2}}}),\frac{1_{F_{Q}}}{\mu(B_{Q})^{\frac{1}{2}}}\rangle|.$
Letting
$e_{Q}:=\frac{1_{E_{Q}}}{\mu(B_{Q})^{\frac{1}{2}}},\qquad h_{Q}:=\frac{1_{F_{Q}}}{\mu(B_{Q})^{\frac{1}{2}}},$
the estimate of $m_{b}^{\phi}$ can be rewritten as
$m_{b}^{\phi}(B_{Q})\lesssim|\langle[b,T](e_{Q}),h_{Q}\rangle|,\quad\mathrm{with}\quad|e_{Q}|+|h_{Q}|\lesssim\frac{1_{B_{Q}^{*}}}{\mu(B_{Q})^{\frac{1}{2}}}.$
For $q\in[1,\infty]$, $p\in(1,\infty)$, $\varepsilon\in(0,1)$ and $\alpha\in(0,\infty)$, we obtain, from the estimates right above followed by an application of Proposition 7.2, that
$\|b\|_{\mathrm{Osc}_{\phi}^{p,q}}(\mu)$ $=\|\{m_{b}^{\phi}(B_{Q})\}_{Q\in\mathscr{D}}\|_{\ell^{p,q}}$
$\lesssim\big\|\big\{|\langle[b,T](e_{Q}),h_{Q}\rangle|\big\}_{Q\in\mathscr{D}}\big\|_{\ell^{p,q}}$
$\lesssim\|\mathcal{M}\|_{L^{2}\times L^{2}\rightarrow L^{1}}\|[b,T]\|_{S^{p,q}}\lesssim\|[b,T]\|_{S^{p,q}}.$
as desired. ∎
8. Characterisation of constants
Motivated by the result from [25, Proposition 4.1], we give a characterisation of constants via Besov-type conditions, with the assumption of the Poincaré inequality.
Proposition 8.1 ([25], Proposition 4.1).
Let $p\in[1,\infty)$ and $(X,\rho,\mu)$ be a doubling metric measure space supporting the $(1,p)$-Poincaré inequality. If $b\in L_{\mathrm{loc}}^{1}(\mu)$ satisfies
(8.1) $(x,y)\mapsto\frac{|b(x)-b(y)|^{p}}{\rho(x,y)^{p}}\frac{1}{V(x,y)}]\in L_{\mathrm{loc}}^{1}(\mu\times\mu),$
then $f$ is equal to a constant almost everywhere.
Corollary 8.2.
Let $d\in[1,\infty)$, $\varepsilon\in[0,1-\frac{1}{d}]$, and $\alpha=\varepsilon d\in[0,d-1]$. Let $(X,\rho,\mu)$ be a space of homogeneous type with lower dimension $d$ and supporting the $(1,p)$-Poincaré inequality with $p=\frac{d}{1+\varepsilon d}=\frac{d}{1+\alpha}$. Let $\phi$ be as in (2.6) and (5.2) If $b\in L_{\mathrm{loc}}^{1}(\mu)$ satisfies
(8.2) $(x,y)\mapsto\frac{|b(x)-b(y)|^{p}}{V(x,y)^{2}}\phi(x,y)^{p}]\in L_{\mathrm{loc}}^{1}(\mu\times\mu),$
then $b$ is equal to a constant almost everywhere. In particular,
(8.3) $\mathbb{B}_{\frac{d}{1+\varepsilon d}}(\phi,\mu)=\dot{B}^{\frac{1}{d}}_{\frac{d}{1+\varepsilon d}}(\mu)=\widetilde{B}^{\alpha}_{\frac{d}{1+\alpha}}(\mu)=\{constants\}.$
Proof..
Suppose that $b$ satisfies (8.2).
Fix a ball $B_{0}:=B(x_{0},R)$. For all $x,y\in B_{0}$, we have
$V(x,y)\lesssim\left(\frac{\rho(x,y)}{R}\right)^{d}\mu(B(x,R))\lesssim\left(\frac{\rho(x,y)}{R}\right)^{d}\mu(B_{0})\lesssim_{B_{0}}\rho(x,y)^{d}.$
Case 1: $\phi(x,y)=V(x,y)^{\varepsilon}$. Then
$\frac{1}{\rho(x,y)^{p}V(x,y)}\lesssim\frac{1}{V(x,y)^{\frac{p}{d}+1}}=\frac{V(x,y)^{p(\frac{1}{p}-\frac{1}{d})}}{V(x,y)^{2}}=\frac{V(x,y)^{p\varepsilon}}{V(x,y)^{2}}=\frac{\phi(x,y)^{p}}{V(x,y)^{2}}.$
Case 2: $\phi(x,y)=\rho(x,y)^{\alpha}$. Then
$\frac{1}{\rho(x,y)^{p}V(x,y)}=\frac{\rho(x,y)^{\alpha p}}{\rho(x,y)^{(1+\alpha)p}V(x,y)}=\frac{\rho(x,y)^{\alpha p}}{\rho(x,y)^{d}V(x,y)}\lesssim\frac{\rho(x,y)^{\alpha p}}{V(x,y)^{2}}=\frac{\phi(x,y)^{p}}{V(x,y)^{2}}.$
Hence, in either case,
$\frac{|b(x)-b(y)|^{p}}{\rho(x,y)^{p}}\frac{1}{V(x,y)}\lesssim\frac{|b(x)-b(y)|^{p}}{V(x,y)^{2}}\phi(x,y)^{p},$
which is integrable over $B_{0}\times B_{0}$ by (8.2). Since this holds for every ball $B_{0}$, we see that (8.1) is satisfied. Then Proposition 8.1 shows that $b$ is constant.
Clearly, every $b\in\mathbb{B}_{p}(\phi,\mu)$ satisfies (8.2), and hence is constant by what we just proved. It follows from the definitions that
$\mathbb{B}_{p}(V^{\varepsilon},\mu)=\dot{B}_{p}^{\frac{1}{p}-\varepsilon}(\mu)=\dot{B}_{\frac{d}{1+\varepsilon d}}^{\frac{1}{d}}(\mu),\qquad\mathbb{B}_{p}(\rho^{\alpha},\mu)=\widetilde{B}_{p}^{\alpha}(\mu)=\widetilde{B}^{\alpha}_{\frac{d}{1+\alpha}}(\mu),$
and hence both these spaces consists only of constants. ∎
9. Proof of the main results
Proof of Theorem 2.14.
The conclusion of Theorem 2.14 (1) is directly deduced from Corollary 4.2. By combination of Proposition 5.2 (equivalence of the Besov norm and the oscillation norm) and Proposition 7.1 (domination of the oscillation norm by the Schatten norm), we obtain the result of Theorem 2.14 (2). Finally, Theorem 2.14 (3) follows from Corollary 8.2. ∎
Proof of Corollary 2.15 (1):.
Since $p\geq 2$, $\varepsilon\in(0,1)$ and $\alpha\in(0,\infty)$, applying Theorem 2.14 (1), we get
$\left\|[b,T_{\varepsilon}]\right\|_{S^{p}}\lesssim\left\|b\right\|_{\dot{B}_{p}^{\frac{1}{p}-\varepsilon}(\mu)}\qquad\mathrm{and}\qquad\left\|[b,\widetilde{T_{\alpha}}]\right\|_{S^{p}}\lesssim\left\|b\right\|_{\widetilde{B}_{p}^{\alpha}(\mu)}.$
Since $p>1$ and $\varepsilon\in(0,\frac{1}{p})\subset(0,1)$, applying Theorem 2.14 (2), we get
$\left\|b\right\|_{\dot{B}_{p}^{\frac{1}{p}-\varepsilon}(\mu)}\lesssim\left\|[b,T_{\varepsilon}]\right\|_{S^{p}}\qquad\mathrm{and}\qquad\left\|b\right\|_{\widetilde{B}_{p}^{\alpha}(\mu)}\lesssim\left\|[b,\widetilde{T_{\alpha}}]\right\|_{S^{p}}.$
Combining the above two inequalities, we complete the proof of Corollary 2.15 (1). ∎
Proof of Corollary 2.15 (2):.
Since $p\in(1,2)\subset(1,\infty)$ and $\varepsilon\in(\max\{0,\frac{1}{p}-\frac{1}{d}\},\frac{1}{p})\subset(0,\frac{1}{p})$, applying 2.14 (2), we obtain
$\left\|b\right\|_{\dot{B}_{p}^{\frac{1}{p}-\varepsilon}(\mu)}\lesssim\left\|[b,T_{\varepsilon}]\right\|_{S^{p}}\qquad\mathrm{and}\qquad\left\|b\right\|_{\widetilde{B}_{p}^{\alpha}(\mu)}\lesssim\left\|[b,\widetilde{T_{\alpha}}]\right\|_{S^{p}},$
which complete the proof of Corollary 2.15 (2). ∎
Proof of Corollary 2.15 (3).
$\Leftarrow$”: It is obvious that if $b$ is constants, then $b,T_{\varepsilon}],[b,\widetilde{T_{\alpha}}]\in S^{p}$ for all relevant parameters values.
$\Rightarrow$”: For $\varepsilon\in(0,\frac{1}{p}-\frac{1}{d}]\cap(0,1-\frac{1}{d})$, we have $p\leq\frac{d}{1+d\varepsilon}=:q$, where $\frac{1}{q}=\frac{1}{d}+\varepsilon<1$ and hence $q>1$. We apply Theorem 2.14 (2) to $q$ in place of $p$ and (2.20) to get
$[b,T_{\varepsilon}]\in S^{p}\subset S^{q}=S^{\frac{d}{1+d\varepsilon}}\,\Rightarrow\,b\in\dot{B}_{\frac{d}{1+d\varepsilon}}^{\frac{1}{d}}(\mu)\equiv\{constants\},$
and
$[b,\widetilde{T_{\alpha}}]\in S^{p}\subset S^{q}=S^{\frac{d}{1+\alpha}}\,\Rightarrow\,b\in\widetilde{B}_{\frac{d}{1+\alpha}}^{\alpha}(\mu)\equiv\{constants\}.$
Proof of Corollary 1.1.
Observe that for $0<\varepsilon<\frac{1}{2}-\frac{1}{d}$, we have $\frac{d}{1+d\varepsilon}>2$. By Remark 2.13, the space $(X,\rho,\mu)$ also satisfies the $(1,\frac{d}{1+d\varepsilon})$-Poincaré inequality.
Part (i): If $p\in[\frac{d}{1+d\varepsilon},\infty)\subset(2,\infty)$, then by Theorem 2.14 (1) and (2), we obtain
(9.1) $\begin{cases}[b,T_{\varepsilon}]\in S^{p}&\iff\,\,\,b\in\dot{B}_{p}^{\frac{1}{p}-\varepsilon}(\mu),\\ [b,\widetilde{T_{\alpha}}]\in S^{p}&\iff\,\,\,b\in\widetilde{B}_{p}^{\alpha}(\mu).\end{cases}$
In particular, this holds for $p\in(\frac{d}{1+d\varepsilon},\infty)$.
Part (ii): If $0<p\leq\frac{d}{1+d\varepsilon}=:q$, then by taking $p=q$ in (9.1) and Theorem 2.14 (3), one has
$[b,T_{\varepsilon}]\in S^{p}\subset S^{q}=S^{\frac{d}{1+d\varepsilon}}\,\Rightarrow\,b\in\dot{B}_{\frac{d}{1+d\varepsilon}}^{\frac{1}{d}}(\mu)\equiv\{constants\},$
and
$[b,\widetilde{T_{\alpha}}]\in S^{p}\subset S^{q}=S^{\frac{d}{1+\alpha}}\,\Rightarrow\,b\in\widetilde{B}_{\frac{d}{1+\alpha}}^{\alpha}(\mu)\equiv\{constants\}.$
10. Fractional integrals with regular kernels
Our main results in this paper, whose proofs we have just completed, only depend on the size and not on the any regularity of the kernel $K$ of the fractional integral. Nevertheless, many examples of fractional integrals arising in applications also possess additional regularity, and we hence some comments on this situation. In particular, we observe that, with some mild regularity of the kernel, the strong non-degeneracy that we have assumed (a uniform estimate over all points in certain balls) already follows from a simpler non-degeneracy condition involving a pair of points only.
Definition 10.1.
A $\phi$-fractional integral kernel (Definition 2.2) is called
  1. (1)
    $\omega$-regular if it satisfies
    (10.1) $|K(x,y)-K(x^{\prime},y)|+|K(y,x)-K(y,x^{\prime})|\leq\omega\Big(\frac{\rho(x,x^{\prime})}{\rho(x,y)}\Big)\frac{\phi(x,y)}{V(x,y)}$
    for $\rho(x,x^{\prime})<(2A_{0})^{-1}\rho(x,y)$, where $\omega$ is a bounded function with $\lim_{t\to 0}\omega(t)=0$;
  2. (2)
    non-degenerate if there are positive constants $c_{0}$ and $\overline{C}$ such that for every $x\in X$ and $r>0$, there exists a point $y\in B(x,\overline{C}r)\setminus B(x,r)$ such that
    (10.2) $|K(x,y)|+|K(y,x)|\geq c_{0}\cdot\frac{\phi(x,y)}{V(x,y)},$
The following lemma clarifies the connection between the two versions of non-degeneracy.
Lemma 10.2.
Let $K$ be a $\phi$-fractional integral kernel (Definition 2.2). If $K$ is non-degenerate and $\omega$-regular (Definition 10.1), then $K$ is strongly non-degenerate (Definition 2.3).
Proof..
Given $x_{0}\in X$ and $r>0$, we apply the condition of non-degeneracy with $x_{0}$ in place of $x$ and $Ar$ in place of $r$, where $A$ is yet to be chosen. This gives us a point $y_{0}\in B(x_{0},\overline{C}Ar)\setminus B(x_{0},Ar)$ such that
$\left|K(x_{0},y_{0})\right|+\left|K(y_{0},x_{0})\right|\geq c_{0}\frac{\phi(x_{0},y_{0})}{V(x_{0},y_{0})},$
thus
$\left|K(x_{0},y_{0})\right|\geq\frac{c_{0}}{2}\frac{\phi(x_{0},y_{0})}{V(x_{0},y_{0})},\quad\text{or}\quad\left|K(y_{0},x_{0})\right|\geq\frac{c_{0}}{2}\frac{\phi(x_{0},y_{0})}{V(x_{0},y_{0})}.$
We assume the first case and proceed to prove (2.8). In the second case, we obtain (2.9) analogously. Since $y_{0}\in B(x_{0},\overline{C}Ar)\setminus B(x_{0},Ar)$, it follows that $\phi(x_{0},y_{0})\sim\phi(x_{0},Ar)$ and $V(x_{0},y_{0})\sim V(x_{0},Ar)$.
Let $x\in B(x_{0},r)$, $y\in B(y_{0},r)$. When $A$ is large, it is clear that we can apply the $\omega$-regularity below to estimate
(10.3) $\begin{split}\left|K(x,y)-K(x_{0},y_{0})\right|&\leq\left|K(x,y)-K(x_{0},y)\right|+\left|K(x_{0},y)-K(x_{0},y_{0})\right|\\ &\leq\omega\Big(\frac{\rho(x,x_{0})}{\rho(x_{0},y)}\Big)\frac{\phi(x_{0},y)}{V(x_{0},y)}+\omega\Big(\frac{\rho(y,y_{0})}{\rho(x_{0},y_{0})}\Big)\frac{\phi(x_{0},y_{0})}{V(x_{0},y_{0})}\\ &\lesssim\Big[\omega\Big(\frac{r}{\rho(x_{0},y)}\Big)+\omega\Big(\frac{r}{\rho(x_{0},y_{0})}\Big)\Big]\frac{\phi(x_{0},y_{0})}{V(x_{0},y_{0})}.\end{split}$
Here $Ar\leq\rho(x_{0},y_{0})\leq A_{0}(\rho(x_{0},y)+\rho(y,y_{0}))<A_{0}\rho(x_{0},y)+A_{0}r$, and hence
$\rho(x_{0},y)>\frac{A-A_{0}}{A_{0}}r\geq 2A_{0}r$
provided that $A\geq A_{0}+2A_{0}^{2}$.
Substituting back to (10.3), we obtain
$\left|K(x,y)-K(x_{0},y_{0})\right|\lesssim\Big[\omega\Big(\frac{A_{0}}{A-A_{0}}\Big)+\omega\Big(\frac{1}{A}\Big)\Big]\frac{\phi(x_{0},y_{0})}{V(x_{0},y_{0})}.$
Recalling that $\lim_{t\to 0}\omega(t)=0$, choosing $A=A(\varepsilon)$ large enough, we can guarantee that
$\left|K(x,y)-K(x_{0},y_{0})\right|\leq\varepsilon\frac{\phi(x_{0},y_{0})}{V(x_{0},y_{0})}$
for any given $\varepsilon>0$. Then clearly $\left|K(x,y)\right|\geq\left|K(x_{0},y_{0})\right|-\left|K(x,y)-K(x_{0},y_{0})\right|$ has the required lower bound.
Moreover, denoting by $v$ the complex unit in the direction of $K(x,y)$, we have
$\bar{v}K(x,y)=\bar{v}K(x_{0},y_{0})+\bar{v}[K(x,y)-K(x_{0},y_{0})]=:\sigma+\zeta=\sigma(1+\frac{\zeta}{\sigma}),$
where $\sigma=\left|K(x_{0},y_{0})\right|\gtrsim\phi(x_{0},y_{0})V(x_{0},y_{0})^{-1}=:\tau$, and $\left|\zeta\right|\leq\varepsilon\tau$; hence $\left|\zeta/\sigma\right|\lesssim\varepsilon$. Then
$\left|\arg(\bar{v}K(x,y))\right|=\left|\arg(1+\frac{\zeta}{\sigma})\right|\lesssim\varepsilon$
can also made as small as desired with sufficiently small $\varepsilon$, i.e., sufficiently large $A$.
Noting that, for $A$ large but fixed,
$\frac{\phi(x_{0},y_{0})}{V(x_{0},y_{0})}\sim\frac{\phi(x_{0},Ar)}{V(x_{0},Ar)}\sim\frac{\phi(x_{0},r)}{V(x_{0},r)},$
this completes the proof. ∎
Example 10.3.
The basic fractional integrals $I_{\varepsilon}$ and $\tilde{I}_{\alpha}$ from (1.4) are strongly non-degenerate $\phi$-fractional integrals with $\phi(x,y)=V(x,y)^{\varepsilon}$ and $\phi(x,y)=\rho(x,y)^{\alpha}$, respectively, but not necessarily $\omega$-regular.
Proof..
The fact that $I_{\varepsilon}$ and $\tilde{I}_{\alpha}$ are strongly non-degenerate $\phi$-fractional integrals was already observed in Example 2.7.
In a general space of homogeneous type, $V(x,y)$ need not be continuous, and hence $\omega$-regularity may easily fail. For example, let $n\geq 2$ and
$X=\{x=(x_{i})_{i=1}^{n}\in\mathbb{R}^{n}\mid\exists i:x_{i}\in\mathbb{Z}\}$
with the $\ell^{\infty}$ metric and the (n1)(n-1)-dimensional Lebesgue measure. This is an Ahlors (n1)(n-1)-regular space of homogeneous type, but V(x,y)V(x,y)ăhas jumps at the points where $x$ or $y$ has more than one integer coordinate. ∎
Remark 10.4.
By Lemma 10.2, $\omega$-regularity together with non-degeneracy is a sufficient condition for strong non-degeneracy, and this is often convenient in applications, but Example 2.7 shows that this condition is not necessary; instead, strong non-degeneracy without any regularity is a strictly more general property. Hence we have formulated the main theorem below in terms of this latter condition.
11. Fractional integrals arising from heat kernels
There is an extensive literature on heat kernels on manifolds and more general metric measure spaces, see e.g. [13, 17, 29]. In this section, we show that negative fractional powers of the generator of a semigroup with a heat kernel are metric fractional integrals in the sense of this paper. In particular, their commutators will be in the scope of our results. We will detail this application in the specific setting of fractional Bessel operators in Section 12, but we first deal with a more general setting here.
We begin by giving the definition of heat kernels and recalling related facts from [17].
Definition 11.1 (Heat kernel).
Let $(X,\rho,\mu)$ be a metric measure space. A family $\{p_{t}\}_{t>0}$ of measurable functions on $X\times X$ is called a heat kernel if for all $s,t>0$ and almost all $x,y\in X$, it satisfies
(i) $p_{t}(x,y)\geq 0;$
(ii) $\int_{X}p_{t}(x,y)\,d\mu(y)=1;$
(iii) $p_{t}(x,y)=p_{t}(y,x);$
(iv) $p_{s+t}(x,y)=\int_{X}p_{s}(x,z)p_{t}(z,y)\,d\mu(z);$
(v) $\lim_{t\to 0^{+}}\int_{X}p_{t}(x,y)f(y)\,d\mu(y)=f(x)\quad\text{in the $L^{2}$-sense for all }f\in L^{2}(X,\mu).$
For any $t>0$, the heat kernel $p_{t}(\cdot,\cdot)$ is the kernel of an operator $P_{t}$, which we write as $e^{-t\mathcal{L}}$. The generator $\mathcal{L}$ of the semigroup $\{e^{-t\mathcal{L}}\}_{t>0}$ is defined by
$\mathcal{L}f:=\lim_{t\to 0}\frac{f-e^{-t\mathcal{L}}(f)}{t},$
for those $f\in L^{2}(X,\mu)$ for which the limit exists in $L^{2}(X,\mu)$. The operator $\mathcal{L}$ is self-adjoint and positive definite. For any $s>0$, the corresponding fractional operator $\mathcal{L}^{-s}$ is defined by
(11.1) $\mathcal{L}^{-s}f(x)=\frac{1}{\Gamma(s)}\int_{0}^{\infty}e^{-t\mathcal{L}}(f)(x)\frac{dt}{t^{1-s}},\quad\forall x\in X.$
Next, we consider the following assumptions for the heat kernel $p_{t}(\cdot,\cdot)$. Let $\gamma>0$. Suppose that $\Phi_{i}:[0,\infty)\rightarrow[0,\infty)$ is a non-negative function for any $i=1,2,3$. The heat kernel $p_{t}(\cdot,\cdot)$ satisfies the following conditions:
  1. (i)
    For all $x,y\in X$,
    (11.2) $p_{t}(x,y)\lesssim\frac{1}{V(x,t^{\gamma})}\Phi_{1}\big(\frac{\rho(x,y)}{t^{\gamma}}\big);$
  2. (ii)
    For all $x,y\in X$,
    (11.3) $p_{t}(x,y)\gtrsim\frac{1}{V(x,t^{\gamma})}\Phi_{2}\big(\frac{\rho(x,y)}{t^{\gamma}}\big);$
  3. (iii)
    For $\rho(x,x^{\prime})<(2A_{0})^{-1}\rho(x,y)$ and some $\varepsilon>0$,
    (11.4) $|p_{t}(x,y)-p_{t}(x^{\prime},y)|+|p_{t}(y,x)-p_{t}(y,x^{\prime})|\lesssim\frac{1}{V(x,t^{\gamma})}\big(\frac{\rho(x,x^{\prime})}{t^{\gamma}}\big)^{\varepsilon}\Phi_{3}\big(\frac{\rho(x,y)}{t^{\gamma}}\big).$
Since the fundamental results of Li and Yau [29] proving bounds of this type for the heat kernel of the Laplace–Beltrami operator on a complete Riemannian manifold of nonnegative Ricci curvature (in which case $\gamma=\frac{1}{2}$ and $\Phi_{j}(u)=e^{-c_{j}u^{2}}$), many further situations giving rise to such heat kernel bounds have been explored in the literature. A very general form of such bounds has been recently studied in [13].
The following is the main result of this section:
Proposition 11.2.
Let $\gamma,s,\varepsilon>0$. Let $(X,\rho,\mu)$ be a space of homogeneous type with a lower dimension $d>0$ and an upper dimension $D>0$. Suppose that the fractional operator $\mathcal{L}^{-s}$ is defined as in (11.1) and $p_{t}(\cdot,\cdot)$ is the heat kernel associated to $\mathcal{L}$. Let $K_{s}(\cdot,\cdot)$ be the kernel of the fractional operator $\mathcal{L}^{-s}$. Then the following statements hold:
  1. (i)
    If $p_{t}(\cdot,\cdot)$ satisfies (11.2) and $\Phi_{1}$ satisfies
    (11.5) $\int_{0}^{1}\Phi_{1}(w)\,w^{d-\frac{s}{\gamma}}\frac{dw}{w}+\int_{1}^{\infty}\Phi_{1}(w)\,w^{D-\frac{s}{\gamma}}\frac{dw}{w}<\infty,$
    then for all $x,y\in X$,
    (11.6) $K_{s}(x,y)\lesssim\frac{\rho(x,y)^{\frac{s}{\gamma}}}{V(x,y)};$
  2. (ii)
    If $p_{t}(\cdot,\cdot)$ satisfies (11.3) and $\Phi_{2}$ is non-zero in a set of positive measure, then for all $x,y\in X$,
    (11.7) $K_{s}(x,y)\gtrsim\frac{\rho(x,y)^{\frac{s}{\gamma}}}{V(x,y)};$
  3. (iii)
    If $p_{t}(\cdot,\cdot)$ satisfies (11.4) and the function $\Phi_{3}$ satisfies
    (11.8) $\int_{0}^{1}\Phi_{3}(w)\,w^{d+\varepsilon-\frac{s}{\gamma}}\frac{dw}{w}+\int_{1}^{\infty}\Phi_{3}(w)\,w^{D+\varepsilon-\frac{s}{\gamma}}\frac{dw}{w}<\infty,$
    then for $\rho(x,x^{\prime})<(2A_{0})^{-1}\rho(x,y)$,
    (11.9) $|K_{s}(x,y)-K_{s}(x^{\prime},y)|+|K_{s}(y,x)-K_{s}(y,x^{\prime})|\lesssim\frac{\rho(x,y)^{\frac{s}{\gamma}}}{V(x,y)}\big(\frac{\rho(x,x^{\prime})}{\rho(x,y)}\big)^{\varepsilon}.$
Proof..
By the doubling and reverse doubling property of the measure $V$, we consider the following two cases. For $0<u\leq\rho(x,y)$,
(11.10) $\big(\frac{\rho(x,y)}{u}\big)^{d}\lesssim\frac{V(x,y)}{V(x,u)}=\frac{V(x,\rho(x,y))}{V(x,u)}\lesssim\big(\frac{\rho(x,y)}{u}\big)^{D},$
and for $u\geq\rho(x,y)$,
(11.11) $\big(\frac{\rho(x,y)}{u}\big)^{D}\lesssim\frac{V(x,y)}{V(x,u)}=\frac{V(x,\rho(x,y))}{V(x,u)}\lesssim\big(\frac{\rho(x,y)}{u}\big)^{d}.$
By (11.4) and a change of variable $u=t^{\gamma}$, we have for all $x,y\in X$,
(11.12) $K_{s}(x,y)$ $=\frac{1}{\Gamma(s)}\int_{0}^{\infty}p_{t}(x,y)\frac{dt}{t^{1-s}}=\frac{1}{\gamma\cdot\Gamma(s)}\int_{0}^{\infty}p_{u^{1/\gamma}}(x,y)\frac{du}{u^{1-\frac{s}{\gamma}}}$
$\sim\Big(\int_{0}^{\rho(x,y)}+\int_{\rho(x,y)}^{\infty}\Big)p_{u^{1/\gamma}}(x,y)\frac{du}{u^{1-\frac{s}{\gamma}}}$
$=:I+II,$
and for $\rho(x,x^{\prime})<(2A_{0})^{-1}\rho(x,y)$,
(11.13) $|K_{s}(x,y)-K_{s}(x^{\prime},y)|+|K_{s}(y,x)-K_{s}(y,x^{\prime})|$
$\lesssim\int_{0}^{\infty}|p_{t}(x,y)-p_{t}(x^{\prime},y)|+|p_{t}(y,x)-p_{t}(y,x^{\prime})|\frac{dt}{t^{1-s}}$
$\lesssim\rho(x,x^{\prime})^{\varepsilon}\int_{0}^{\infty}\frac{1}{V(x,t^{\gamma})}\Phi_{3}\big(\frac{\rho(x,y)}{t^{\gamma}}\big)\frac{dt}{t^{1+\varepsilon\gamma-s}}$
$\sim\rho(x,x^{\prime})^{\varepsilon}\int_{0}^{\infty}\frac{1}{V(x,u)}\Phi_{3}\big(\frac{\rho(x,y)}{u}\big)\frac{du}{u^{1+\varepsilon-\frac{s}{\gamma}}}$
$=\rho(x,x^{\prime})^{\varepsilon}\Big(\int_{0}^{\rho(x,y)}+\int_{\rho(x,y)}^{\infty}\Big)\frac{1}{V(x,u)}\Phi_{3}\big(\frac{\rho(x,y)}{u}\big)\frac{du}{u^{1+\varepsilon-\frac{s}{\gamma}}}$
$=:III+IV.$
To proceed, we estimate the above terms II-IVIV.
First,
[…]
$=\frac{\rho(x,y)^{D}}{V(x,y)}\int_{1}^{\infty}\Phi_{1}(w)\,\bigl(w^{-1}\rho(x,y)\bigr)^{\frac{s}{\gamma}-D}\,\frac{dw}{w}\qquad\quad\text{by }w=\frac{\rho(x,y)}{u}$
[…]
Similarly,
[…]
$=\frac{\rho(x,y)^{d}}{V(x,y)}\int_{0}^{1}\Phi_{1}(w)\,\bigl(w^{-1}\rho(x,y)\bigr)^{\frac{s}{\gamma}-d}\,\frac{dw}{w}\qquad\quad\text{by }w=\frac{\rho(x,y)}{u}$
[…]
and
(11.16) $III$ $\lesssim\frac{\rho(x,y)^{\frac{s}{\gamma}}}{V(x,y)}\big(\frac{\rho(x,x^{\prime})}{\rho(x,y)}\big)^{\varepsilon}\int_{1}^{\infty}\Phi_{3}(w)w^{D+\varepsilon-\frac{s}{\gamma}}\,\frac{dw}{w}\lesssim\frac{\rho(x,y)^{\frac{s}{\gamma}}}{V(x,y)}\big(\frac{\rho(x,x^{\prime})}{\rho(x,y)}\big)^{\varepsilon},$
and
(11.17) $IV$ $\lesssim\frac{\rho(x,y)^{\frac{s}{\gamma}}}{V(x,y)}\big(\frac{\rho(x,x^{\prime})}{\rho(x,y)}\big)^{\varepsilon}\int_{0}^{1}\Phi_{3}(w)w^{d+\varepsilon-\frac{s}{\gamma}}\,\frac{dw}{w}\lesssim\frac{\rho(x,y)^{\frac{s}{\gamma}}}{V(x,y)}\big(\frac{\rho(x,x^{\prime})}{\rho(x,y)}\big)^{\varepsilon},$
by (11.8) in the last steps of both (11.16) and (11.17). This completes the proof of (11.6) and (11.9).
It remains to consider the lower bound in (11.7). First,
[…]
$=\frac{\rho(x,y)^{d}}{V(x,y)}\int_{1}^{\infty}\Phi_{2}(w)\,\bigl(w^{-1}\rho(x,y)\bigr)^{\frac{s}{\gamma}-d}\,\frac{dw}{w}\qquad\text{by }w=\frac{\rho(x,y)}{u}$
$=\frac{\rho(x,y)^{\frac{s}{\gamma}}}{V(x,y)}\int_{1}^{\infty}\Phi_{2}(w)\,w^{d-\frac{s}{\gamma}}\,\frac{dw}{w}\gtrsim\frac{\rho(x,y)^{\frac{s}{\gamma}}}{V(x,y)},$
assuming, in the last step, that $\Phi_{2}$ is non-zero in a subset of $1,\infty)$ of positive measure.
Similarly,
[…]
$=\frac{\rho(x,y)^{D}}{V(x,y)}\int_{0}^{1}\Phi_{2}(w)\,\bigl(w^{-1}\rho(x,y)\bigr)^{\frac{s}{\gamma}-D}\,\frac{dw}{w}\qquad\text{by }w=\frac{\rho(x,y)}{u}$
$=\frac{\rho(x,y)^{\frac{s}{\gamma}}}{V(x,y)}\int_{0}^{1}\Phi_{2}(w)\,w^{D-\frac{s}{\gamma}}\,\frac{dw}{w}\gtrsim\frac{\rho(x,y)^{\frac{s}{\gamma}}}{V(x,y)},$
assuming, in the last step, that $\Phi_{2}$ is non-zero in a subset of $[0,1]$ of positive measure.
Since $\Phi_{2}$ is non-zero in a subset of $0,\infty)$ of positive measure, at least one the two lower bounds above is valid. Summing up, and noting that both terms are certainly non-negative, it follows that
$K_{s}(x,y)=I+II\gtrsim\frac{\rho(x,y)^{\frac{s}{\gamma}}}{V(x,y)},$
which complete the proof of (11.7). ∎
Remark 11.3.
We now present some examples of functions to illustrate Proposition 11.2. These examples are used to provide intuition for the proposition and will be applied in the next section. If $\Phi_{1}$ is bounded on $[0,1]$, then for any $0<s<\gamma d$,
$\int_{0}^{1}\Phi_{1}(w)\,w^{d-\frac{s}{\gamma}}\frac{dw}{w}\lesssim\int_{0}^{1}w^{d-\frac{s}{\gamma}}\frac{dw}{w}\sim 1,$
If the function $u^{a}\,\Phi_{1}(u)$ is bounded on $1,\infty)$ for some $a>D-\frac{s}{\gamma}$, then
$\int_{1}^{\infty}\Phi_{1}(w)\,w^{D-\frac{s}{\gamma}}\frac{dw}{w}=\int_{1}^{\infty}\Phi_{1}(w)w^{a}\cdot w^{D-a-\frac{s}{\gamma}}\frac{dw}{w}\lesssim\int_{1}^{\infty}w^{D-a-\frac{s}{\gamma}}\frac{dw}{w}\sim 1.$
Similarly, if the function $\Phi_{3}$ is bounded on $[0,1]$ and $u^{\widetilde{a}}\,\Phi_{3}(u)=0$ is bounded on $1,\infty)$ for some $\widetilde{a}>D+\varepsilon-\frac{s}{\gamma}$, then for any $0<s<\gamma(d+\varepsilon)$, the function $\Phi_{3}$ satisfies (11.8).
In particular, if $\Phi_{j}(u)=\exp(-c_{j}u^{d_{j}})$ with $c_{j},d_{j}>0$, then the conclusions of Proposition 11.2 are valid for all $s\in(0,\gamma d)$.
12. Application to fractional Bessel operators
In this section, we apply our main results to the fractional Bessel operator. We first recall the (non-fractional) Bessel operator from Huber [20]. For $n\geq 0$ and $\lambda>0$, the $(n+1)$-dimension Bessel operator $\Delta_{\lambda}^{(n+1)}$ on $\mathbb{R}^{n+1}_{+}:=\mathbb{R}^{n}\times(0,\infty)$ is defined by (1.5). When $n=0$, we write $\mathbb{R}_{+}:=(0,\infty)$. The operator $\Delta_{\lambda}^{(n+1)}$ is symmetric and non-negative in $L^{2}(\mathbb{R}^{n+1}_{+},dm_{\lambda}^{{(n+1)}})$, where
$dm_{\lambda}^{{(n+1)}}(x):=x_{n+1}^{2\lambda}dx.$
For simplicity, we write $\Delta_{\lambda}:=\Delta_{\lambda}^{(n+1)}$ and $m_{\lambda}:=m_{\lambda}^{{(n+1)}}$ when the dimension is clear from the context.
In [3], Betancor et al. gave the kernel estimates of the one-dimensional fractional Bessel operator $(-\Delta_{\lambda}^{(1)})^{-\alpha/2}$ on $(0,\infty)$ with parameters $n=0$, $\lambda>0$ and $0<\alpha<1+2\lambda$. In this section, we extend their result to the corresponding fractional Bessel operator $(-\Delta_{\lambda}^{(n+1)})^{-\alpha/2}$ for all $n\geq 0$, $\lambda>0$ and $0<\alpha<n+1+2\lambda$. Taking $\mathcal{L}=-\Delta_{\lambda}$ and $s=\alpha/2$ in (11.1), the fractional Bessel operator is defined by
(12.1) $(-\Delta_{\lambda})^{-\alpha/2}f(x)=\frac{1}{\Gamma(\alpha/2)}\int_{0}^{\infty}e^{t\Delta_{\lambda}}(f)(x)\frac{dt}{t^{1-\alpha/2}},\quad\forall x\in\mathbb{R}^{n+1}_{+}.$
Corollary 12.1.
Let $n\geq 0$, $\lambda>0$ and $0<\alpha<n+1$. Let $(-\Delta_{\lambda})^{-\alpha/2}$ be the fractional Bessel operator in $L^{2}(\mathbb{R}^{n+1}_{+},\left|x-y\right|,m_{\lambda})$. Let $S^{p}:=S^{p}(L^{2}(dm_{\lambda}^{(n+1)}))$ and $\widetilde{B}_{p}(m_{\lambda}):=\widetilde{B}_{p}^{\alpha}(m_{\lambda}^{(n+1)})$ be defined as in (2.16) with $(X,\rho,\mu)=(\mathbb{R}^{n+1}_{+},\left|x-y\right|,m_{\lambda})$. Then the following conclusions hold for all $b\in L_{\mathrm{loc}}^{1}(\mathbb{R}^{n+1}_{+})$:
  1. (1)
    If $p\in[2,\infty)$ and $\alpha\in(0,\frac{n+1}{p})$, then $b,(-\Delta_{\lambda})^{-\alpha/2}]\in S^{p}$ if and only if $b\in\widetilde{B}_{p}^{\alpha}(m_{\lambda})$.
  2. (2)
    If $p\in(1,2)$ and $\alpha\in(\frac{n+1}{p}-1,\frac{n+1}{p})$, then $b,(-\Delta_{\lambda})^{-\alpha/2}]\in S^{p}$ only if $b\in\widetilde{B}_{p}^{\alpha}(m_{\lambda})$.
  3. (3)
    If $p\in(0,n+1)$ and $\alpha\in(0,n)\cap(0,\frac{n+1}{p}-1]$, then $b,(-\Delta_{\lambda})^{-\alpha/2}]\in S^{p}$ if and only if $b$ is constant.
We intend to apply Corollary 2.15 to prove Corollary 12.1. (Similarly, Corollary 1.1 implies Corollary 1.3, as we already sketched in the Introduction.) Therefore, we need to verify that the space $(X,\rho,\mu)=(\mathbb{R}_{+}^{n+1},|x-y|,m_{\lambda})$ and the operator $\tilde{T}_{\alpha}=(-\Delta_{\lambda})^{-\frac{\alpha}{2}}$ satisfy the assumptions of Corollary 2.15. Concerning the space, we can quote the following results:
Lemma 12.2 ([11], Eq. (2.2)).
Let $n\geq 0$ and $\lambda>0$. For any $x\in\mathbb{R}^{n+1}_{+}$ and $r>0$, let $B_{\mathbb{R}^{n+1}_{+}}(x,r)=B(x,r)\,\cap\mathbb{R}^{n+1}_{+}$. Then for every $x=(x_{1},\dots,x_{n+1})\in\mathbb{R}^{n+1}_{+}$ and $r>0$,
(12.2) $V_{\lambda}(x,r):=m_{\lambda}(B_{\mathbb{R}^{n+1}_{+}}(x,r))\sim r^{n+1}x_{n+1}^{2\lambda}+r^{n+1+2\lambda}.$
Proposition 12.3 ([23], Proposition 4.2).
For every $n\geq 0$ and $\lambda>0$, the space $(\mathbb{R}^{n+1}_{+},|x-y|,m_{\lambda})$ is a space of homogeneous type with lower dimension $d=n+1$ and upper dimension $D=n+1+2\lambda$. Moreover, $(\mathbb{R}^{n+1}_{+},|x-y|,m_{\lambda})$ satisfies the $(1,1)$-Poincaré inequality.
By Remark 2.13, $(\mathbb{R}^{n+1}_{+},|x-y|,m_{\lambda})$ also satisfies the $(1,\frac{n+1}{\alpha+1})$-Poincaré inequality for all $\alpha\in(0,n)$.
It remains to show the relevant conditions for the kernel of the fractional Bessel operators. We will derive these from bounds for the heat kernels associated to Bessel operators from [10, Section 7.5].
  1. (1)
    If $n=0$, then $\Delta_{\lambda}=\Delta_{\lambda}^{(1)}$ and the heat kernel associated to $\Delta_{\lambda}^{(1)}$ is
    (12.3) $W_{t}^{\lambda}(x,y)=\frac{(xy)^{-\lambda+\frac{1}{2}}}{2t}e^{-\frac{x^{2}+y^{2}}{4t}}I_{\lambda-\frac{1}{2}}(\frac{xy}{2t}),\quad\forall x,y\in\mathbb{R}_{+},$
    where $I_{\nu}$ is the modified Bessel function of the first kind with order $\nu>-1$ (see [4] for more details). The kernel of the operator $(\Delta_{\lambda}^{(1)})^{-\alpha/2}$ is obtained by
    (12.4) $K_{\lambda,\alpha}(x,y)=\frac{1}{\Gamma(\alpha/2)}\int_{0}^{\infty}W_{t}^{\lambda}(x,y)\frac{dt}{t^{1-\alpha/2}}.$
    for any $x,y\in\mathbb{R}_{+}$ and $x\neq y$.
  2. (2)
    If $n\geq 1$, then the operator $\Delta_{\lambda}=\Delta_{\lambda}^{(n+1)}$ can be written as $\Delta_{\lambda}=\Delta^{(n)}+\Delta_{\lambda}^{(1)}$, where $\Delta^{(n)}$ denotes the standard Laplacian on $\mathbb{R}^{n}$ and $\Delta_{\lambda}^{(1)}$ denotes the Bessel operator on $\mathbb{R}_{+}$. Hence, it is clear that $e^{-t\Delta_{\lambda}}=e^{-t\Delta^{(n)}}\cdot e^{-t\Delta_{\lambda}^{(1)}}$ and the heat kernel associated to $\Delta_{\lambda}$ is
    (12.5) $K_{e^{-t\Delta_{\lambda}}}(x,y)=K_{e^{-t\Delta^{(n)}}}(x^{\prime},y^{\prime})\cdot W_{t}^{\lambda}(x_{n+1},y_{n+1}),$
    for $x=(x^{\prime},x_{n+1}),y=(y^{\prime},y_{n+1})\in\mathbb{R}^{n}\times(0,\infty)$, where K_e^-tΔ^(n)(x’,y’):=e^-|x’-y’|24t (4πt)^-n/2, is the heat kernel of standard Laplacian and $W_{t}^{\lambda}$ is the heat kernel of $\Delta_{\lambda}^{(1)}$ as in (12.3). The kernel of the operator $(\Delta_{\lambda}^{(n+1)})^{-\alpha/2}$ is obtained by (12.1),
    (12.6) $K_{\lambda,\alpha}(x,y)=\frac{1}{\Gamma(\alpha/2)}\int_{0}^{\infty}K_{e^{-t\Delta^{(n)}}}(x^{\prime},y^{\prime})\cdot W_{t}^{\lambda}(x_{n+1},y_{n+1})\frac{dt}{t^{1-\alpha/2}}.$
    for any $x=(x^{\prime},x_{n+1}),y=(y^{\prime},y_{n+1})\in\mathbb{R}^{n}\times(0,\infty)$ and $x\neq y$.
Letting $n=0$ in (12.6), we get (12.4) by interpreting $K_{e^{-t\Delta^{(0)}}}\equiv 1$. Hence, (12.6) is well defined for all $n\geq 0$.
Remark 12.4.
Since $K_{e^{-t\Delta^{(n)}}}(x^{\prime},y^{\prime})=K_{e^{-t\Delta^{(n)}}}(y^{\prime},x^{\prime})$ and, by (12.3), also
$W_{t}^{\lambda}(x_{n+1},y_{n+1})=W_{t}^{\lambda}(y_{n+1},x_{n+1}),$
it is immediate that $K_{\lambda,\alpha}(x,y)=K_{\lambda,\alpha}(y,x)$ for any $n\geq 0$, $x,y\in\mathbb{R}^{n+1}_{+}$ with $x\neq y$.
As for the case $n=0$, the size estimate and the smoothness estimate of the kernel $K_{{\lambda,\alpha}}$ associated with the fractional Bessel operator have been proved by Betancor et.al. in [3, Lemmas 5.1, 5.2]. The following Proposition 12.5 extends their result to all $n\geq 0$ and gives the lower bound of the kernel $K_{{\lambda,\alpha}}$.
Proposition 12.5.
Let $n\geq 0$, $\lambda>0$ and $0<\alpha<n+1$. Then for all $x,y\in\mathbb{R}_{+}^{n+1}$ with $x\neq y$,
(12.7) $K_{\lambda,\alpha}(x,y)\sim\frac{|x-y|^{\alpha}}{V_{\lambda}(x,y)}:=\frac{|x-y|^{\alpha}}{m_{\lambda}(B_{\mathbb{R}^{n+1}_{+}}(x,|x-y|))},$
and for $|x-x^{\prime}|<|x-y|/2$,
(12.8) $|K_{\lambda,\alpha}(x,y)-K_{\lambda,\alpha}(x^{\prime},y)|+|K_{\lambda,\alpha}(y,x)-K_{\lambda,\alpha}(y,x^{\prime})|\lesssim\frac{|x-y|^{\alpha}}{V_{\lambda}(x,y)}\cdot\frac{|x-x^{\prime}|}{|x-y|}.$
In particular, the fractional Bessel operator $(-\Delta_{\lambda})^{-\alpha/2}$ is a strongly non-degenerate metric fractional integral operator with kernel $K_{\lambda,\alpha}$ satisfying (2.7), (10.1), and the non-degenerate condition (10.2), where $\phi(x,y)=|x-y|^{\alpha}$, $V(x,y)=V_{\lambda}(x,y)$, and $\omega(t)=t$.
We will derive Proposition 12.5 from related heat kernel bounds using Proposition 11.2, which holds in particular for $(X,\rho,\mu)=(\mathbb{R}^{n+1}_{+},|x-y|,m_{\lambda})$. This requires checking the relevant assumptions for the heat kernel associated to Bessel operators $\Delta_{\lambda}$. The relevant upper bounds have already been verified in [11]:
Lemma 12.6 ([11], Lemma 2.4).
Let $n\geq 0$. For all multi-indices $\beta\in\mathbb{N}^{n+1}$, there exists positive constants $C_{\beta},c>0$ such that for all $x,y\in\mathbb{R}_{+}^{n+1}$,
(12.9) $|\partial_{x}^{\beta}K_{e^{-t\Delta_{\lambda}}}(x,y)|+|\partial_{y}^{\beta}K_{e^{-t\Delta_{\lambda}}}(x,y)|\leq\frac{C_{\beta}}{t^{\frac{|\beta|}{2}}V_{\lambda}(x,\sqrt{t})}\exp(-c\frac{|x-y|^{2}}{t}).$
To obtain the required lower bound, we first present the following lemma on the modified Bessel functions $I_{\nu}$:
Lemma 12.7.
For any $u>0$ and $\nu>-1$, the modified Bessel functions satisfy
(12.10) $I_{\nu}(u)\sim\begin{cases}u^{\nu}&\textup{if }u\in(0,1],\\ u^{-\frac{1}{2}}e^{u}&\textup{if }u\in[1,\infty).\end{cases}$
Proof..
We recall the asymptotic properties of the modified Bessel function $I_{\nu}$ from [4], p.109. For any $\nu>-1$,
$\lim_{u\rightarrow 0^{+}}u^{-\nu}I_{\nu}(u)=\frac{1}{2^{\nu}\Gamma(\nu+1)},$
and
$\lim_{u\rightarrow+\infty}\frac{\sqrt{2\pi u}}{e^{u}}\cdot I_{\nu}(u)=1.$
This implies that there exist constants $0<\varepsilon<1$ and $M>1$ such that
$I_{\nu}(u)\sim\begin{cases}u^{\nu}&\textup{if }u\in(0,\varepsilon],\\ u^{-\frac{1}{2}}e^{u}&\textup{if }u\in[M,\infty).\end{cases}$
On the other hand, one of definitions of $I_{\nu}(u)$ for any $u>0$ and $\nu>-1$ is
$I_{\nu}(u)=\sum_{n=0}^{\infty}\frac{(u/2)^{2n+\nu}}{\Gamma(n+1)\Gamma(n+\nu+1)}.$
This shows that the function $I_{\nu}$ is positive and continuous. Consequently,
$I_{\nu}(u)\sim 1\sim u^{\nu}\sim u^{-\frac{1}{2}}e^{u}\quad\quad\textup{if }u\in[\varepsilon,M].$
Hence, we get (12.10). ∎
We establish the following lower bound of the heat kernel $K_{e^{-t\Delta_{\lambda}}}$.
Lemma 12.8.
Let $n\geq 0$ and $\delta>0$. For all $x,y\in\mathbb{R}_{+}^{n+1}$,
(12.11) $K_{e^{-t\Delta_{\lambda}}}(x,y)\gtrsim\frac{1}{V_{\lambda}(x,\sqrt{t})}\exp\Big({-(\frac{1}{4}+\delta)\cdot\frac{|x-y|^{2}}{t}}\Big).$
Proof..
We now take $r=\sqrt{t}$ in (12.2), which implies that
(12.12) $V_{\lambda}(x,\sqrt{t})\sim t^{\frac{n+1}{2}}(x_{n+1}^{2\lambda}+t^{\lambda}).$
By using the identity
$e^{-\frac{|x-y|^{2}}{4t}}=e^{-\frac{|x^{\prime}-y^{\prime}|^{2}}{4t}}\cdot e^{-\frac{|x_{n+1}-y_{n+1}|^{2}}{4t}},$
we rewrite the heat kernel (12.5) (where $W_{t}^{\lambda}$ is given in (12.3)) as
(12.13) $K_{e^{-t\Delta_{\lambda}}}(x,y)$ $=\frac{1}{(4\pi t)^{\frac{n}{2}}}e^{-\frac{|x^{\prime}-y^{\prime}|^{2}}{4t}}\frac{(x_{n+1}y_{n+1})^{-\lambda+\frac{1}{2}}}{2t}I_{\lambda-\frac{1}{2}}(\frac{x_{n+1}y_{n+1}}{2t})e^{-\frac{x_{n+1}^{2}+y_{n+1}^{2}}{4t}}$
$\sim t^{-\frac{n+1+2\lambda}{2}}e^{-\frac{|x-y|^{2}}{4t}}\cdot e^{-\frac{x_{n+1}y_{n+1}}{2t}}\big(\frac{x_{n+1}y_{n+1}}{2t}\big)^{-\lambda+\frac{1}{2}}I_{\lambda-\frac{1}{2}}(\frac{x_{n+1}y_{n+1}}{2t})$
$=:t^{-\frac{n+1+2\lambda}{2}}e^{-\frac{|x-y|^{2}}{4t}}D_{t}(x_{n+1},y_{n+1}).$
Case 1: $\frac{x_{n+1}y_{n+1}}{2t}<1$. Taking $\nu=\lambda-\frac{1}{2}$ in (12.10), we conclude that
(12.14) $D_{t}(x_{n+1},y_{n+1})$ $\sim e^{-\frac{x_{n+1}y_{n+1}}{2t}}\big(\frac{x_{n+1}y_{n+1}}{2t}\big)^{-\lambda+\frac{1}{2}}\,\big(\frac{x_{n+1}y_{n+1}}{2t}\big)^{\lambda-\frac{1}{2}}$
$=e^{-\frac{x_{n+1}y_{n+1}}{2t}}\sim 1.$
Hence
[…]
$\geq\frac{1}{V_{\lambda}(x,\sqrt{t})}e^{-(\frac{1}{4}+\delta)\cdot\frac{|x-y|^{2}}{t}}.$
Case 2: $\frac{x_{n+1}y_{n+1}}{2t}>1$. We note that
$x_{n+1}y_{n+1}$ $\leq x_{n+1}(x_{n+1}+|x_{n+1}-y_{n+1}|)=x_{n+1}^{2}+x_{n+1}|x_{n+1}-y_{n+1}|$
$\leq 2x_{n+1}^{2}+|x_{n+1}-y_{n+1}|^{2}\lesssim x_{n+1}^{2}+|x-y|^{2}.$
Hence
(12.16) $(x_{n+1}y_{n+1})^{\lambda}$ $\lesssim x_{n+1}^{2\lambda}+|x-y|^{2\lambda}=x_{n+1}^{2\lambda}+t^{\lambda}\frac{|x-y|^{2\lambda}}{t^{\lambda}}\lesssim(x_{n+1}^{2\lambda}+t^{\lambda})e^{\delta\frac{|x-y|^{2}}{t}},$
using $1\leq e^{x}$ and $x^{a}\lesssim e^{x}$ for all $x,a>0$ in the last inequality.
Taking $\nu=\lambda-\frac{1}{2}$ in (12.10) and recalling that $\frac{x_{n+1}y_{n+1}}{2t}>1$, we conclude that
(12.17) $D_{t}(x_{n+1},y_{n+1})$ $\sim e^{-\frac{x_{n+1}y_{n+1}}{2t}}\big(\frac{x_{n+1}y_{n+1}}{2t}\big)^{-\lambda+\frac{1}{2}}\cdot e^{\frac{x_{n+1}y_{n+1}}{2t}}\big(\frac{x_{n+1}y_{n+1}}{2t}\big)^{-\frac{1}{2}}$
$=\big(\frac{x_{n+1}y_{n+1}}{2t}\big)^{-\lambda}\gtrsim t^{\lambda}(x_{n+1}^{2\lambda}+t^{\lambda})^{-1}e^{-\delta\frac{\left|x-y\right|^{2}}{t}}.$
Thus we obtain
[…]
Hence, combining (12.15) with (12.18), we obtain (12.11). ∎
Proof of Proposition 12.5: Choose parameters
$s=\frac{\alpha}{2},\,\gamma=\frac{1}{2},\,d=n+1,\,\varepsilon=1,\,D=n+1+2\lambda$
and
$\mathcal{L}=-\Delta_{\lambda},\,p_{t}=K_{e^{-t\Delta_{\lambda}}},\,K_{s}=K_{\lambda,\alpha},\,V(\cdot,\cdot):=V_{\lambda}(\cdot,\cdot)$
in Proposition 11.2. Taking $|\beta|=0,1$ in Lemma 12.6 and by Lemma 12.8, we obtain that $K_{e^{-t\Delta_{\lambda}}}$ satisfies (11.2)-(11.4) with $\Phi_{j}(u)=e^{-c_{j}u^{2}}$, where $c_{1}=c_{3}=c$ is the same as in Lemma 12.6, and $c_{2}=\frac{1}{4}+\delta$ for some $\delta>0$.
Obviously, $\Phi_{2}$ is positive in a set of positive measure. By Remark 11.3, it is easy to verify that the functions $\Phi_{1},\Phi_{3}$ satisfy conditions (11.5) and (11.8), respectively. Then, by Proposition 11.2, the kernel $K_{s}=K_{\lambda,\alpha}$ satisfies (11.6), (11.7) and (11.9). Thus, we obtain the desired estimates (12.7) and (12.8).